### v2

# Exact overlaps for ”all” integrable matrix product states of rational spin chains

Abstract
The overlaps between integrable matrix product states (MPS) and Bethe states are important in both the non-equilibrium statistical physics and the AdS/CFT duality. We present the general MPS overlap formula. The result is a product of a ratio of Gaudin determinants and a prefactor. The Gaudin determinants depend on the spin chain but not on the MPS. The MPS dependent prefactor is given for all integrable MPS of the $\mathfrak{gl}_{N}$, $\mathfrak{o}_{N}$ and $\mathfrak{sp}_{N}$ symmetric spin chains with arbitrary representations.
I Introduction
Over the past ten years, intensive research has been carried out on the topic of integrable boundary states. The overlaps between boundary and Bethe states have important applications in two areas of theoretical physics: non-equilibrium dynamics in statistical physics and the AdS/CFT duality.
A major breakthrough in the last decade has been the laboratory simulation of integrable statistical physics models [1, 2, 3]. A common feature is the experimental study of so-called quantum quenches, where certain initial states (ground state of the pre-quenched system) are prepared and the resulting finite-time dynamics are monitored. This launched an intensive theoretical investigation of the non-equilibrium dynamics of integrable models [4, 5]. These studies are based on the so-called Quench Action method, which allows the calculation of the long-term equilibrium state and the possibility to study the time evolution of specific initial states [6, 7, 8]. The Quench Action method is based on explicit knowledge of the overlaps between the initial state and the Bethe eigenstates. This method has been successfully applied to a number of different models [9, 10, 11, 12], but the investigated initial states were the so-called two-site product states which have essentially minimal entanglement.
In parallel, the investigation of the AdS/CFT duality with defects by integrability methods was started. First, it was found that the one-point functions can be traced back to overlaps of certain matrix product states (MPS) and Bethe states for the so-called D3-D5 type domain wall defect of the $\mathcal{N}=4$ super Yang-Mills theory ($\mathcal{N}=4$ SYM) [13, 14]. These overlap formulas, which give the one-point functions at weak coupling, have been conjectured and verified numerically [15]. Later, boundary states at finite coupling were defined [16, 17], and their corresponding overlap functions give the finite coupling one-point functions in the asymptotic limit [18, 19]. This domain wall setup was also generalized for the ABJM theory [20, 21]. Similar boundary states have been found for other defects. In the presence of giant graviton some correlation functions are also given by boundary state overlaps in the $\mathcal{N}=4$ SYM [22] and in the ABJM [23]. Boundary state overlaps also appear for the one-point functions in the presence of Wilson- and ’t Hooft lines [24, 25]. Most recently, the integrable boundary states and their overlaps emerged in the investigation of the Coulomb branch of $\mathcal{N}=4$ SYM [26]. The integrable conformal defects in $\mathcal{N}=4$ SYM was also classified [27]. The physical significance of these defects is that they violate conformal [26] or super-symmetry [28] in certain configurations, which can lead to more realistic theories. In these models, the role of the boundary state overlaps is important.
Due to the above statistical physical and AdS/CFT duality applications, the research on the overlaps of integrable boundary and Bethe states has been initiated. The definition of the integrable boundary states appeared first in the context of 2d quantum field theories [29] and later was generalized to spin chains [30, 31]. The first important observation on the overlaps was that they are proportional to the ratio of Gaudin determinants [32]. The Gaudin determinants are universal, i.e., they do not depend on the boundary state. The boundary state dependent part of the overlap function is the prefactor before the ratio. Later this observation was used to conjecture overlap formulas [33, 15, 34]. There have been only a few proofs [32, 35, 36], mainly concentrated on spin 1/2 XXX or XXZ spin chains. It was not clear how these methods could be generalized to general spins or higher rank symmetries. Later, a method based on the algebraic Bethe Ansatz was developed to determine and prove the overlap formulas of $\mathfrak{gl}_{N}$ symmetric spin chains for any representations and any two-site product states [37, 38, 39]. Although these results are extremely general, they do not say anything about MPS overlaps. In terms of the applications, the MPS overlaps are particularly important. For quantum quenches, the ground state of the pre-quench system can be approximated better with an MPS than with a two-site product state. Even in the AdS/CFT duality, the most interesting defects correspond to MPS boundary states.
In this letter we present the general MPS overlap formula (21). Every integrable MPS corresponds to a $K$-matrix which is a representation of a reflection algebra. One can define a series of commuting matrices ($F$-operators) from the components of the $K$-matrix. The spectrum of the $F$-operators completely defines the MPS overlap formula. The definition of the $F$-operators depends only on the reflection algebra. In the Supplemental Materials we give the definitions of the $F$-operators for every reflection algebras of the $\mathfrak{gl}_{N}$, $\mathfrak{o}_{N}$ or $\mathfrak{sp}_{N}$ symmetric spin chains. These results give the overlaps for all integrable MPS of the $\mathfrak{gl}_{N}$, $\mathfrak{o}_{N}$ or $\mathfrak{sp}_{N}$ symmetric spin chains with arbitrary representations.
II Integrable spin chains and matrix product states
II.1 Definitions for integrable spin chains
In this letter we do not focus on a particular model but we work with a wide class of integrable spin chains. A spin chain is given by a Hamiltonian $H$. The $H$ acts on a Hilbert space $\mathcal{H}$ with a tensor product form $\mathcal{H}=\left(\mathbb{C}^{d}\right)^{J}$ where $d$ is the dimension of the one-site Hilbert space and $J$ is the length of the spin chain. The model is integrable if there exists a generating function $\mathcal{T}(u)$ of the commuting charges which commute with the Hamiltonian
$\left[H,\mathcal{T}(u)\right]=0.$ (1)
The generating function $\mathcal{T}(u)$ is the transfer matrix. Let us introduce an auxiliary space $\mathbb{C}^{N}$. For integrable spin chains with periodic boundary condition the transfer matrix is
$\mathcal{T}(u)=\sum_{j=1}^{N}T_{j,j}(u),$ (2)
where $T_{i,j}(u)$ are the entries of the monodromy matrix for $i,j=1,\dots,N$. We also introduce the matrix notation in the auxiliary space $T(u)=\sum_{i,j}e_{i,j}\otimes T_{i,j}(u)\in\mathrm{End}(\mathbb{C}^{N}\otimes\mathcal{H})$, where $e_{i,j}$ are the unit matrices of $\mathbb{C}^{N}$ which have the components $\left(e_{i,j}\right)_{a,b}=\delta_{i,a}\delta_{j,b}$. The monodromy matrix satisfies the famous $RTT$-relation
$R_{1,2}(u-v)T_{1}(u)T_{2}(v)=T_{2}(v)T_{1}(u)R_{1,2}(u-v),$ (3)
where the $R(u)\in\mathrm{End}(\mathbb{C}^{N}\otimes\mathbb{C}^{N})$ is the R-matrix and the subscripts $1,2$ label two copies of the auxiliary space. The $RTT$-relation has two consequences: the R-matrix has to satisfy the famous Yang-Baxter (YB) equation
$R_{1,2}(u-v)R_{1,3}(u)R_{2,3}(v)=R_{2,3}(v)R_{1,3}(u)R_{1,2}(u-v),$ (4)
and the transfer matrix generates commuting operators
$\left[\mathcal{T}(u),\mathcal{T}(v)\right]=0.$ (5)
In this letter we investigate rational spin chains which have Lie-algebra symmetry $\mathfrak{gl}_{N}$, $\mathfrak{o}_{N}$ or $\mathfrak{sp}_{2n}$. The $\mathfrak{gl}(N)$ symmetric R-matrix is
$R(u)=\mathbf{1}+\frac{1}{u}\mathbf{P},\quad\mathbf{P}=\sum_{i,j=1}^{N}e_{i,j}\otimes e_{j,i}.$ (6)
The $\mathfrak{o}_{N}$ and $\mathfrak{sp}_{N}$ symmetric R-matrices can be written as
$R(u)=\mathbf{1}+\frac{1}{u}\mathbf{P}-\frac{1}{u+\kappa_{N}}\mathbf{Q},\quad\mathbf{Q}=\sum_{i,j=1}^{N}\theta_{i}\theta_{j}e_{i,j}\otimes e_{\bar{i},\bar{j}},$ (7)
where $\bar{i}=N+1-i$ and $\kappa_{N}=\frac{N\mp 2}{2}$. The minus or plus signs correspond to the $\mathfrak{o}_{N}$ or $\mathfrak{sp}_{N}$ algebras, respectively. For $\mathfrak{o}_{N}$ $\theta_{i}=+1$ and for $\mathfrak{sp}_{N}$
$\theta_{i}=\begin{cases}+1,&i=1,\dots,\frac{N}{2},\\ -1,&i=\frac{N}{2}+1,\dots,N.\end{cases}$ (8)
The monodromy matrices are representations of the Yangian algebras $Y(\mathfrak{g})$ [40], where $\mathfrak{g}=\mathfrak{gl}_{N},\mathfrak{o}_{N}$ or $\mathfrak{sp}_{2n}$.
Let us introduce the eigenvalues and eigenvectors of the transfer matrix
$\mathcal{T}(u)|\bar{u}\rangle=\tau(u|\bar{u})|\bar{u}\rangle,$ (9)
where $|\bar{u}\rangle$ is the Bethe vector and $\bar{u}\equiv\left\{\bar{u}^{j}\right\}_{j=1}^{n}$ denotes the set of Bethe roots $\bar{u}^{j}=\left\{u_{k}^{j}\right\}_{k=1}^{r_{j}}$. The $n$ is the rank of the symmetry algebra ($n$ for $\mathfrak{gl}_{n+1}$, $\mathfrak{o}_{2n+1}$, $\mathfrak{sp}_{2n}$ and $\mathfrak{o}_{2n}$). The Bethe roots should satisfy the Bethe Ansatz equations [41]
$\alpha_{a}(u_{k}^{a})\prod_{b=1}^{n}\prod_{l=1}^{r_{b}}\frac{u_{k}^{a}-u_{l}^{b}-\frac{i}{2}C_{a,b}}{u_{k}^{a}-u_{l}^{b}+\frac{i}{2}C_{a,b}}=-1,$ (10)
for $a=1,\dots,n$ and $k=1,\dots,r_{a}$. The $C_{a,b}$ are the inner products of the simple roots of the corresponding algebra  [^1] We give the explicit form in the Supplemental Materials., the functions $\alpha_{a}(u)$ depends on the representations of the Hilbert space and the inhomogeneities. A set of the Bethe roots $\bar{u}^{a}$ corresponds to a node $a$ of the Dynkin diagram for $a=1,\dots,n$, see figure 1.
Figure 1: Dynkin diagrams (first row $\mathfrak{gl}_{n+1}$, second row $\mathfrak{o}_{2n+1}$ or $\mathfrak{sp}_{2n}$, third row $\mathfrak{o}_{2n}$) and the non-trivial involutions which correspond to achiral pair structures.
II.2 Integrable matrix product states
For even $J$ let us define the following matrix product state (MPS)
$\langle\mathrm{MPS}|=\\ \sum_{i_{1},\dots,i_{J}}\mathrm{Tr}\left[\psi_{i_{J-1},i_{J}}^{(J/2)}\dots\psi_{i_{1},i_{2}}^{(1)}\right]\langle i_{1},i_{2},\dots,i_{J-1},i_{J}|,$ (11)
where $\psi_{i,j}^{(k)}\in\mathrm{End}(\mathbb{C}^{d_{B}})$ for $i,j=1,\dots,d$, $k=1,\dots,J/2$ and $d_{B}$ is the boundary (or bond) dimension. The $\psi_{i,j}^{(k)}$ matrices can be different (i.e. $\psi_{i,j}^{(k)}\neq\psi_{i,j}^{(l)}$) for inhomogeneous chains but they are given by the same functions as $\psi_{i,j}^{(k)}\equiv\psi_{i,j}(\theta_{k})$, where $\theta_{k}$-s are the inhomogeneities. For homogenous chains it can happen that $\psi_{i,j}^{(k)}=\omega_{j}\omega_{i}$ for which the MPS can be also defined for even or odd length:
$\langle\mathrm{MPS}|=\sum_{i_{1},\dots,i_{J}}\mathrm{Tr}\left[\omega_{i_{J}}\dots\omega_{i_{1}}\right]\langle i_{1},i_{2},\dots,i_{J-1},i_{J}|.$ (12)
We also define states by ”cutting the trace”
$\langle\Psi_{\alpha,\beta}|=\\ \sum_{i_{1},\dots,i_{J}}\left(\psi_{i_{J-1},i_{J}}^{(J/2)}\dots\psi_{i_{1},i_{2}}^{(1)}\right)_{\alpha,\beta}\langle i_{1},i_{2},\dots,i_{J-1},i_{J}|,$ (13)
where $\alpha,\beta=1,\dots,d_{B}$ are indices of the boundary space. The boundary state is integrable if it satisfies the $KT$-relation
$\sum_{k,\gamma}K_{i,k}^{\alpha,\gamma}(u)\langle\Psi_{\gamma,\beta}|T_{k,j}(u)=\sum_{k,\gamma}\langle\Psi_{\alpha,\gamma}|\widehat{T}_{i,k}(-u)K_{k,j}^{\gamma,\beta}(u),$ (14)
where the coefficients $K_{i,j}^{\alpha,\beta}(u)\in\mathbb{C}$ define the K-matrix $\mathbf{K}_{i,j}(u)=\sum_{\alpha,\beta}K_{i,j}^{\alpha,\beta}(u)e_{\alpha,\beta}^{B}\in\mathrm{End}(\mathbb{C}^{d_{B}})$, $\mathbf{K}(u)=\sum_{i,j}e_{i,j}\otimes\mathbf{K}_{i,j}(u)\in\mathrm{End}(\mathbb{C}^{N}\otimes\mathbb{C}^{d_{B}})$. We used the elementary matrices $(e_{\alpha,\beta}^{B})_{a,b}=\delta_{\alpha,a}\delta_{\beta,b}$ of the boundary space $\mathbb{C}^{d_{B}}$. We also define the notation $\langle\Psi|\in\mathcal{H}^{*}\otimes\mathrm{End}(\mathbb{C}^{d_{B}})$ which is a covector in the Hilbert space and a matrix in the boundary space: $\langle\Psi|=\sum_{\alpha,\beta}\langle\Psi_{\alpha,\beta}|\otimes e_{\alpha,\beta}^{B}$. Using these matrix notations in the boundary space the $KT$-relation has the following compact form
$\mathbf{K}(u)\langle\Psi|T(u)=\langle\Psi|\widehat{T}(-u)\mathbf{K}(u).$ (15)
Here the tensors $\mathbf{K}(u)$, $\langle\Psi|$ or $T(u)$ act trivially on the Hilbert-, the auxiliary- or the boundary spaces, respectively.
There are two types of $KT$-relation: the uncrossed $KT$-relation where $\widehat{T}=T$ and the crossed $KT$-relation where
$\sum_{k=1}^{N}\widehat{T}_{k,i}(u)T_{k,j}(u)=\lambda_{1}(u)\lambda_{1}(-u)\delta_{i,j}\mathbf{1}.$ (16)
Here we used the pseudo-vacuum eigenvalue $T_{1,1}(u)|0\rangle=\lambda_{1}(u)|0\rangle$  [^2] The pseudo-vacuum is the Bethe vector without Bethe roots, i.e., $r_{1}=r_{2}=\dots=r_{n}=0$.. Assuming that the $K$-marix is invertible, the condition
$\langle\mathrm{MPS}|\mathcal{T}(u)=\langle\mathrm{MPS}|\widehat{\mathcal{T}}(-u),$ (17)
is following from the $KT$-relation.
$\mathfrak{g}$ $\mathfrak{h}$ KT-relation P.s. $n_{+}$
$\mathfrak{gl}_{N}$ $\mathfrak{o}_{N}$ crossed chiral $N-1$
$\mathfrak{sp}_{N}$ crossed chiral $N-1$
$\mathfrak{gl}_{M}\oplus\mathfrak{gl}_{N-M}$ uncrossed achiral $\left\lfloor\frac{N}{2}\right\rfloor$
$\mathfrak{o}_{2n+1}$ $\mathfrak{o}_{M}\oplus\mathfrak{o}_{2n+1-M}$ uncrossed chiral $n$
$\mathfrak{sp}_{2n}$ $\mathfrak{sp}_{2m}\oplus\mathfrak{sp}_{2n-2m}$ uncrossed chiral $n$
$\mathfrak{gl}_{n}$ uncrossed chiral $n$
$\mathfrak{o}_{2n=4k}$ $\mathfrak{o}_{2m}\oplus\mathfrak{o}_{2n-2m}$ uncrossed chiral $n$
$\mathfrak{o}_{2m+1}\oplus\mathfrak{o}_{2n-2m-1}$ uncrossed achiral $n-1$
$\mathfrak{gl}_{n}$ uncrossed chiral $n$
$\mathfrak{o}_{2n=4k+2}$ $\mathfrak{o}_{2m}\oplus\mathfrak{o}_{2n-2m}$ uncrossed achiral $n-1$
$\mathfrak{o}_{2m+1}\oplus\mathfrak{o}_{2n-2m-1}$ uncrossed chiral $n$
$\mathfrak{gl}_{n}$ uncrossed achiral $n-1$
Table 1: The reflections algebras $Y(\mathfrak{g},\mathfrak{h})$ and the corresponding pair structures.
Another consequence of the $KT$-relation is that the $K$-matrix satisfies the reflection equation
$R_{1,2}(u-v)\mathbf{K}_{1}(-u)\bar{R}_{1,2}(u+v)\mathbf{K}_{2}(-v)=\\ \mathbf{K}_{2}(-v)\bar{R}_{1,2}(u+v)\mathbf{K}_{1}(-u)R_{1,2}(u-v),$ (18)
where
$\bar{R}_{1,2}(u)=\begin{cases}R_{1,2}(u),&\text{for the uncrossed }KT,\\ R_{1,2}^{t_{2}}(-u),&\text{for the crossed }KT.\end{cases}$ (19)
The reflection equation defines reflection algebra with the generators $\mathbf{K}(u)$. Let us take the series expansion
$\mathbf{K}(u)=\mathcal{U}\otimes\mathbf{1}+u^{-1}\sum_{i,j=1}^{N}e_{i,j}\otimes\mathbf{k}_{i,j}^{(1)}+\mathcal{O}(u^{-2}),$ (20)
where $\mathcal{U}\in\mathrm{Aut}(\mathbb{C}^{N})$ and $\mathbf{k}_{i,j}^{(1)}\in\mathrm{End}(\mathbb{C}^{d_{B}})$. The matrices $\mathbf{k}_{i,j}^{(1)}$ are generates Lie-subalgebra $\mathfrak{h}$ [44], and the $K$-matrix is a representation of the twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ [45, 46, 47, 48]. The classification of the twisted Yangians is shown in table 1.
III Exact overlaps of integrable MPS
III.1 Pair structures
Now we continue with the overlaps between the integrable MPS-s and the on-shell Bethe states. The integrability condition (17) is equivalent to non-trivial selection rules for the on-shell overlaps, namely, the on-shell overlap is non-vanishing only when the Bethe roots have pair structures, i.e. $\bar{u}^{j}=-\bar{u}^{\rho(j)}$ for $j=1,\dots,n$. The $\rho$ is an involution of the Dynkin-diagram. If $\rho$ is trivial, i.e., $\rho(j)=j$ for every $j$ then we have chiral pair structure. The pair structure is achiral for non-trivial $\rho$. We have two possibilities for non-trivial $\rho$: for $\mathfrak{gl}_{N}$ $\rho(j)=N-j$; and for $\mathfrak{o}_{2n}$ $\rho(n-1)=n$, $\rho(n)=n-1$ and $\rho(j)=j$ for $j<n-1$. For $\mathfrak{o}_{2n+1}$ and $\mathfrak{so}_{2n}$ there is no non-trivial involution therefore only the chiral pair structure is possible. Every twisted Yangian has a specific pair structure which is summarized in table 1. Integrable MPS with a given chirality can only exist for spin chains where the following conditions are satisfied $\alpha_{j}(u)\alpha_{\rho(j)}(-u)=1$ for $j=1,\dots,n$.
Also, a set of Bethe roots $\bar{u}^{j}$ is chiral if the condition $\rho(j)=j$ holds. For the sake of simplicity, we focus on the states where every chiral set has an even number of elements. We cover the odd cases separately in the Supplemental Materials. The pair structures halve the degrees of freedom, i.e. half of the Bethe roots can be expressed from the other half. More precisely, the full set has a decomposition $\bar{u}=\bar{u}^{+}\cup\bar{u}^{-}$ for which $\bar{u}^{+},\bar{u}^{-}$ are disjoint subsets and $\bar{u}^{-}=-\bar{u}^{+}$ in the pair structure limit. Let $n_{+}$ be the number of sets in $\bar{u}^{+}$. The explicit values of $n_{+}$ are listed in table 1.
III.2 Overlap formula
The on-shell overlaps are
$\frac{\langle\mathrm{MPS}|\bar{u}\rangle}{\sqrt{\langle\bar{u}|\bar{u}\rangle}}={\color[rgb]{0,0,1}\left[\sum_{k=1}^{d_{B}}\beta_{k}\prod_{j=1}^{n_{+}}\tilde{\mathcal{F}}_{k}^{(j)}(\bar{u}^{+,j})\right]}{\color[rgb]{1,0,0}\sqrt{\frac{\det G^{+}}{\det G^{-}}}}.$ (21)
The overlap formula has two parts. The ratio of Gaudin determinants (red) does not depend on MPS. They depend only on the Bethe ansatz equations and the pair structure. The Gaudin determinants are defined in the Supplementary material but they are also found in several previous papers [32, 38, 39].
The prefactor (blue) is the MPS dependent part. We introduced shorthand notation ~$\tilde{\mathcal{F}}_{k}^{(j)}(\bar{v})=\prod_{v_{l}\in\bar{v}}\tilde{\mathcal{F}}_{k}^{(j)}(v_{l})$ where ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$-s are single-variable functions for $j=1,\dots,n_{+}$ and $k=1,\dots,d_{B}$. These functions ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$ depend only on the $K$-matrix, but not on the concrete representation of the Hilbert space. For every twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ there is a commuting subalgebra which are generated by $\mathbf{F}^{(j)}(u)$ for $j=1,\dots,n_{+}$. The $F$-operators $\mathbf{F}^{(j)}(u)$ can be expressed by the components of the $K$-matrix. Since these are commuting operators they can be diagonalized simultaneously and the eigenvalues give ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$.
The prefactor also contains the coefficients $\beta_{k}\in\mathbb{C}$ for $k=1,\dots,d_{B}$. They are the eigenvalues of the pseudo-vacuum overlap operator $\mathbf{B}=\langle\Psi|0\rangle$ which commutes with the $F$-operators.
III.3 Example $Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$
Here we give the concrete functions ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$ for the reflection algebra $Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$. The non-vanishing overlaps require chiral pair structure, i.e., there exist decompositions $\bar{u}^{s}=\bar{u}^{+,s}\cup\bar{u}^{-,s}$ for $s=1,\dots,N-1$. In the pair structure limit $\bar{u}^{-,s}=-\bar{u}^{+,s}$. For chiral pair structure $n_{+}=N-1$ therefore we have $N-1$ $F$-operators.
First we define a series of nested $K$-matrices
$\mathbf{K}_{a,b}^{(k+1)}(u)=\mathbf{K}_{a,b}^{(k)}(u)-\mathbf{K}_{a,k}^{(k)}(u)\left[\mathbf{K}_{k,k}^{(k)}(u)\right]^{-1}\mathbf{K}_{k,b}^{(k)}(u),$ (22)
with the initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. The $F$-operators are given by
$\mathbf{F}^{(k)}(u)=\left[\mathbf{K}_{k,k}^{(k)}(u)\right]^{-1}\mathbf{K}_{k+1,k+1}^{(k+1)}(u),$ (23)
for $k=1,\dots,N-1$. They generate a commuting subalgebra and they also commute with the operator $\mathbf{B}=\langle\Psi|0\rangle$, i.e.
$\left[\mathbf{F}^{(k)}(u),\mathbf{F}^{(l)}(v)\right]=\left[\mathbf{F}^{(k)}(u),\mathbf{B}\right]=0,$ (24)
for $k,l=1,\dots,N-1$. They can be diagonalized simultaneously
$\begin{split}\mathbf{F}^{(k)}(u)&=\mathbf{A}\mathrm{diag}(\mathcal{F}_{1}^{(k)}(u),\dots,\mathcal{F}_{d_{B}}^{(k)}(u))\mathbf{A}^{-1},\\ \mathbf{B}&=\mathbf{A}\mathrm{diag}(\beta_{1},\dots,\beta_{d_{B}})\mathbf{A}^{-1},\end{split}$ (25)
where $\mathbf{A}$ is an invertible matrix in the boundary space. Finally, the ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$ functions of the overlap formula (21) are
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\mathcal{F}_{j}^{(k)}(iu-k/2)\sqrt{\frac{u^{2}}{u^{2}+1/4}}.$ (26)
This result gives the overlaps for any representation of the reflection algebra $Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$. The definitions of the $F$-operators depend only on the reflection algebra. In the Supplementary Material we give the concrete definitions of ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$ for all twisted Yangians $Y(\mathfrak{g},\mathfrak{h})$ of table 1 case by case.
IV How to apply the formula
Although the end result is very general, its application is not a trivial task. In this section, we describe the steps for applying the formula (21). A typical case might be that we have an integrable spin chain with an MPS, and we want to decide whether the MPS is integrable, and if so, we want to write up the overlap formula. In this case we need to follow the steps below. We assume that we know the monodromy matrix which satisfies the $RTT$-relation (3) and commutes with the Hamiltonian of our model (1). For the $\mathfrak{gl}_{N}$ symmetric model we may also need the inverse monodromy matrix (16).
1. Deciding on integrability.
We can check the integrability of the MPS with the equation (17). At this step, note that the spectral parameter of the monodromy matrix can be shifted ($T(u-a)$ also satisfies the defining equations (3), (1)) , therefore it can happen that only the shifted equation
$\langle\mathrm{MPS}|\mathcal{T}(u)=\langle\mathrm{MPS}|\widehat{\mathcal{T}}(-u+b),$ (27)
is satisfied in our initial convention. In this case we only need to change our convention as $T(u)\to T(u+b/2)$ to obtain the integrability condition (17). In the $\mathfrak{gl}_{N}$ case, it must also be decided whether the uncrossed or the crossed equation is satisfied (or neither).
2. Finding the K-matrix.
After we verified the integrability of the MPS (and fixed the convention for the spectral parameter), we have to find the $K$-matrix $\mathbf{K}(u)$ by solving the $KT$-relation (17). We can choose as short spin chain as possible. For the MPS-s which are built from two-site operators (11) we can use $J=2$ and for the one-site operator MPS (12) we can use $J=1$. Solving the KT-relation is a simple task since this equation is linear in the components of $\mathbf{K}(u)$.
3. Identify the reflection algebra
Now we have to find the twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ algebra corresponding to our K-matrix. For this we have to identify the Lie-subalgebra $\mathfrak{h}$ from the series expansion of the $K$-matrix (20). The knowledge of twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ already selects the pair structure which defines the Gaudin determinants (see Supplemental Materials).
4. ”Rotations” of the K-matrix. (if necessary)
It is possible that the MPS is in the ”wrong direction” for the overlap formula. It means the following. The formulas of the $F$-operators (22),(23) assumes that some components of the $K$-matrix are invertible. But this may not be true for our particular MPS. However, in all cases there is a $G=\exp(\mathfrak{g})$ transformation that makes the ”desired” elements invertible. The group element $g\in G$ acts on the MPS and the K-matrix in the following way
$\begin{split}\langle\Psi_{\alpha,\beta}^{g}|&=\langle\Psi_{\alpha,\beta}|\Delta(g),\\ \mathbf{K}^{g}(u)&=\left(\hat{g}^{-1}\otimes\mathbf{1}\right)\mathbf{K}(u)\left(g\otimes\mathbf{1}\right),\end{split}$ (28)
where $\hat{g}=g$ except for the crossed $\mathfrak{gl}_{N}$ when $\hat{g}=\left(g^{t}\right)^{-1}$. The $\Delta(g)$ is the usual co-product corresponding to the Hilbert space. The $G$-action preserves the $KT$-relation. Our task is to find a group element $g(\varphi)=\exp(\varphi x)$ ($x\in\mathfrak{g}$,$\varphi\in\mathbb{C}$) for which the definition of the $F$-operators are well defined. The further steps are performed on the ”rotated” K-matrix $\mathbf{K}^{g(\varphi)}$ (if the rotation is necessary), and we can take the $\varphi\to 0$ limit in the final formula.
5. Calculation of the F-operators
Since we know the $K$-matrix and the associated reflection algebra $Y(\mathfrak{g},\mathfrak{h})$, all we have to do is to calculate the $F$-operators based on the section of $Y(\mathfrak{g},\mathfrak{h})$ in the Supplemental Materials. Their eigenvalues define the functions ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$.
V Conclusions and outlook
In this letter we gave the exact overlap functions (21) for every integrable MPS of the $\mathfrak{gl}_{N}$, $\mathfrak{o}_{N}$ and $\mathfrak{sp}_{N}$ symmetric spin chains with arbitrary representations. For every MPS, there is a corresponding $K$-matrix which is a representation of a reflection algebra. The $K$-matrix defines a series of commuting matrices: the $F$-operators. The spectrum of the $F$-operators defines the MPS overlap formula. In the Supplementary Materials we introduced the $F$-operators for the reflections algebras of the $\mathfrak{gl}_{N}$, $\mathfrak{o}_{N}$ and $\mathfrak{sp}_{N}$ symmetric spin chains.
The formulas of this letter have been verified by extensive numerical tests. The development of precise proofs has been postponed to later publications.
An interesting future direction could be the extension of the results to graded spin chains, such as $\mathfrak{gl}(M|N)$ and $\mathfrak{osp}(M|2n)$ symmetric spin chains, or the Hubbard model. The formulas are likely to be generalizable to the trigonometric cases.
There exists another method to calculate MPS overlaps based on dressing formulas [49]. The advantage of dressing formulas is that they can be applied under general twists, but the disadvantage is that they have to be calculated separately for each representation of the reflection algebra, unlike the formulas in this letter, which are true for all representations.
Acknowledgments
The research was supported by the NKFIH grant PD142929 and the János Bolyai Research Scholarship of the Hungarian Academy of Science.
Supplemental Materials: Exact overlaps for ”all” integrable matrix product states of rational spin chains
S-I Gaudin matrices
In this section we define the Gaudin determinants $\det G^{\pm}$ of the MPS overlap formula
$\frac{\langle\mathrm{MPS}|\bar{u}\rangle}{\sqrt{\langle\bar{u}|\bar{u}\rangle}}=\left[\sum_{k=1}^{d_{B}}\beta_{k}\prod_{j=1}^{n_{+}}\tilde{\mathcal{F}}_{k}^{(j)}(\bar{u}^{+,j})\right]\sqrt{\frac{\det G^{+}}{\det G^{-}}}.$ (S-1)
They depend on the Bethe Ansatz equation and the pair structure. In the following we define them case by case.
The Bethe Ansatz equation are the following.
  • For the $\mathfrak{gl}_{n+1}$ symmetric spin chains the Bethe equations are
    $e^{i\phi_{k}^{a}}:=\alpha_{a}(u_{k}^{a})\prod_{l=1}^{r_{a}}\frac{u_{k}^{a}-u_{l}^{a}-i}{u_{k}^{a}-u_{l}^{a}+i}\prod_{l=1}^{r_{a-1}}\frac{u_{k}^{a}-u_{l}^{a-1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a-1}-\frac{i}{2}}\prod_{l=1}^{r_{a+1}}\frac{u_{k}^{a}-u_{l}^{a+1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a+1}-\frac{i}{2}}=-1,\quad a=1,\dots,n.$ (S-2)
  • For the $\mathfrak{o}_{2n+1}$ symmetric spin chains the Bethe equations are
    $\begin{split}e^{i\phi_{k}^{a}}&:=\alpha_{a}(u_{k}^{a})\prod_{l=1}^{r_{a}}\frac{u_{k}^{a}-u_{l}^{a}-i}{u_{k}^{a}-u_{l}^{a}+i}\prod_{l=1}^{r_{a-1}}\frac{u_{k}^{a}-u_{l}^{a-1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a-1}-\frac{i}{2}}\prod_{l=1}^{r_{a+1}}\frac{u_{k}^{a}-u_{l}^{a+1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a+1}-\frac{i}{2}}=-1,\quad a=1,\dots,n-1,\\ e^{i\phi_{k}^{n}}&:=\alpha_{n}(u_{k}^{n})\prod_{l=1}^{r_{n}}\frac{u_{k}^{n}-u_{l}^{n}-i/2}{u_{k}^{n}-u_{l}^{n}+i/2}\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n}-u_{l}^{n-1}+\frac{i}{2}}{u_{k}^{n}-u_{l}^{n-1}-\frac{i}{2}}=-1.\end{split}$ (S-3)
  • For the $\mathfrak{sp}_{2n}$ symmetric spin chains the Bethe equations are
    $\begin{split}e^{i\phi_{k}^{a}}&:=\alpha_{a}(u_{k}^{a})\prod_{l=1}^{r_{a}}\frac{u_{k}^{a}-u_{l}^{a}-i}{u_{k}^{a}-u_{l}^{a}+i}\prod_{l=1}^{r_{a-1}}\frac{u_{k}^{a}-u_{l}^{a-1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a-1}-\frac{i}{2}}\prod_{l=1}^{r_{a+1}}\frac{u_{k}^{a}-u_{l}^{a+1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a+1}-\frac{i}{2}}=-1,\quad a=1,\dots,n-2,\\ e^{i\phi_{k}^{n-1}}&:=\alpha_{n-1}(u_{k}^{n-1})\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n-1}-u_{l}^{n-1}-i}{u_{k}^{n-1}-u_{l}^{n-1}+i}\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n-1}-u_{l}^{n-2}+\frac{i}{2}}{u_{k}^{n-1}-u_{l}^{n-2}-\frac{i}{2}}\prod_{l=1}^{r_{n}}\frac{u_{k}^{n-1}-u_{l}^{n}+i}{u_{k}^{n-1}-u_{l}^{n}-i}=-1,\\ e^{i\phi_{k}^{n}}&:=\alpha_{n}(u_{k}^{n})\prod_{l=1}^{r_{n}}\frac{u_{k}^{n}-u_{l}^{n}-2i}{u_{k}^{n}-u_{l}^{n}+2i}\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n}-u_{l}^{n-1}+i}{u_{k}^{n}-u_{l}^{n-1}-i}=-1.\end{split}$ (S-4)
  • For the $\mathfrak{o}_{2n}$ symmetric spin chains the Bethe equations are
    $\begin{split}e^{i\phi_{k}^{a}}&:=\alpha_{a}(u_{k}^{a})\prod_{l=1}^{r_{a}}\frac{u_{k}^{a}-u_{l}^{a}-i}{u_{k}^{a}-u_{l}^{a}+i}\prod_{l=1}^{r_{a-1}}\frac{u_{k}^{a}-u_{l}^{a-1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a-1}-\frac{i}{2}}\prod_{l=1}^{r_{a+1}}\frac{u_{k}^{a}-u_{l}^{a+1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a+1}-\frac{i}{2}}=-1,\quad a=1,\dots,n-3,\\ e^{i\phi_{k}^{n-2}}&:=\alpha_{n-2}(u_{k}^{n-2})\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n-2}-u_{l}^{n-2}-i}{u_{k}^{n-2}-u_{l}^{n-2}+i}\prod_{l=1}^{r_{n-3}}\frac{u_{k}^{n-2}-u_{l}^{n-3}+\frac{i}{2}}{u_{k}^{n-2}-u_{l}^{n-3}-\frac{i}{2}}\times\\ &\times\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n-2}-u_{l}^{n-1}+\frac{i}{2}}{u_{k}^{n-2}-u_{l}^{n-1}-\frac{i}{2}}\prod_{l=1}^{r_{n}}\frac{u_{k}^{n-2}-u_{l}^{n}+\frac{i}{2}}{u_{k}^{n-2}-u_{l}^{n}-\frac{i}{2}}=-1,\\ e^{i\phi_{k}^{n-1}}&:=\alpha_{n-1}(u_{k}^{n-1})\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n-1}-u_{l}^{n-1}-i}{u_{k}^{n-1}-u_{l}^{n-1}+i}\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n-1}-u_{l}^{n-2}+\frac{i}{2}}{u_{k}^{n-1}-u_{l}^{n-2}-\frac{i}{2}}=-1,\\ e^{i\phi_{k}^{n}}&:=\alpha_{n}(u_{k}^{n})\prod_{l=1}^{r_{n}}\frac{u_{k}^{n}-u_{l}^{n}-i}{u_{k}^{n}-u_{l}^{n}+i}\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n}-u_{l}^{n-2}+\frac{i}{2}}{u_{k}^{n}-u_{l}^{n-2}-\frac{i}{2}}=-1.\end{split}$ (S-5)
The functions $\alpha_{a}(u)$ depends on the representations of the Hilbert space and the inhomogeneities.
For chiral pair structures $\bar{u}^{s}=-\bar{u}^{s}$ for $s=1,\dots,n$. We define the quantum numbers $n_{+}=n$ and $r_{s}^{+}=r_{s}/2$ for $s=1,\dots,n$. We introduce new notations for the Bethe roots $u_{k}^{+,s}\equiv u_{k}^{s}$, $u_{k}^{-,s}\equiv u_{r_{s}+1-k}^{s}$ where $s=1,\dots,n_{+}$ and $k=1,\dots,r_{s}^{+}$. We also define the following sets $\bar{u}^{\pm,s}=\left\{u_{k}^{\pm,s}\right\}_{k=1}^{r_{s}^{+}}$ for $s=1,\dots,n_{+}$ and $\bar{u}^{\pm}=\left\{\bar{u}^{\pm,1},\bar{u}^{\pm,2},\dots,\bar{u}^{\pm,n_{+}}\right\}$. In the pair structure limit $u_{k}^{-,s}=-u_{k}^{+,s}$ for $s=1,\dots,n_{+}$ and $k=1,\dots,r_{s}^{+}$. Chiral integrable MPS exists only for the spin chains where the following conditions are satisfied: $\alpha_{s}(u)\alpha_{s}(-u)=1$ for $s=1,\dots,n$.
The possible achiral pair structures are the following
[…]
  • For $\mathfrak{o}_{2n}$, the achiral pair structure is $\bar{u}^{1}=-\bar{u}^{1}$, $\bar{u}^{2}=-\bar{u}^{2}$, $\dots$,$\bar{u}^{n-2}=-\bar{u}^{n-2}$ and $\bar{u}^{n}=-\bar{u}^{n-1}$. We can see that the nodes $\bar{u}^{s}$ for $s=1,\dots,n-2$ are chiral. We define the quantum numbers $n_{+}=n-1$, $r_{s}^{+}=r_{s}/2$ for $s=1,\dots,n-2$ and $r_{n_{+}}^{+}=r_{n-1}=r_{n}$. We introduce new notations for Bethe roots $u_{k}^{+,s}\equiv u_{k}^{s}$, $u_{k}^{-,s}\equiv u_{r_{s}+1-k}^{s}$ where $s=1,\dots,n_{+}-1$, $k=1,\dots,r_{s}^{+}$ and $u_{k}^{+,n_{+}}\equiv u_{k}^{n-1}$, $u_{k}^{-,n_{+}}\equiv u_{k}^{n}$ where $k=1,\dots,r_{n-1}^{+}$. Achiral integrable MPS-s of $\mathfrak{o}_{2n}$ symmetric spin chains exist only when the following conditions are satisfied: $\alpha_{s}(u)\alpha_{s}(-u)=1$ for $s=1,\dots,n-2$ and $\alpha_{n-1}(u)\alpha_{n}(-u)=1$.
For all chiral or achiral pair structures we also define the following sets $\bar{u}^{\pm,s}=\left\{u_{k}^{\pm,s}\right\}_{k=1}^{r_{s}^{+}}$ where $s=1,\dots,n_{+}$. We also have $\bar{u}^{\pm}=\left\{\bar{u}^{\pm,1},\bar{u}^{\pm,2},\dots,\bar{u}^{\pm,n_{+}}\right\}$. In the pair structure limit $u_{k}^{-,s}=-u_{k}^{+,s}$ for $s=1,\dots,n_{+}$ and $k=1,\dots,r_{s}^{+}$.
The expressions $\phi_{k}^{s}$ are depend on the Bethe roots and they are defined above. These can be grouped according to the pair structure in the same way as the Bethe roots: $\bar{\phi}^{\pm,s}=\left\{\phi_{k}^{\pm,s}\right\}_{k=1}^{r_{s}^{+}}$ where $s=1,\dots,n_{+}$ and $\bar{\phi}^{\pm}=\left\{\bar{\phi}^{\pm,1},\bar{\phi}^{\pm,2},\dots,\bar{\phi}^{\pm,n_{+}}\right\}$. We can define the components of the Gaudin matrix
$G_{k,l}^{\mu,\nu}=\frac{\partial\phi_{k}^{\mu}}{\partial u_{l}^{\nu}}.$ (S-6)
The Gaudin matrices have $\sum_{s=1}^{n}r_{s}$ rows and columns and the rows are labeled by $\nu,l$ and the columns are labeled by $\mu,k$. According to the decompositions $\bar{u}=\bar{u}^{+}\cup\bar{u}^{-}$ and $\bar{\phi}=\bar{\phi}^{+}\cup\bar{\phi}^{-}$ the Gaudin matrix has the following block form
$G=\left(\begin{array}[]{cc}A^{++}&A^{+-}\\ A^{-+}&A^{--}\end{array}\right),$ (S-7)
where
$\left(A^{\sigma_{1}\sigma_{2}}\right)_{k,l}^{\mu,\nu}=\frac{\partial\phi_{k}^{\sigma_{1},\mu}}{\partial u_{l}^{\sigma_{2},\nu}},$ (S-8)
where $\sigma_{1},\sigma_{2}\in\left\{+,-\right\}$. The Gaudin determinants are
$\det G^{\pm}=\det(A^{++}\pm A^{+-}).$ (S-9)
It is crucial that we first take the derivative and only after take the pair structure limit.
S-II Extra selection rules
Besides the pair structure, there may be other selection rules for non-vanishing overlaps. These depend on the Hilbert space. In the Hilbert space there is a special vector, the pseudo-vacuum $|0\rangle$ (Bethe vector without Bethe roots $|0\rangle\equiv|\emptyset\rangle$) and we need the $\mathfrak{g}$ quantum numbers of the pseudo-vacuum. These can be obtained from the series expansion of $\alpha$-functions
$\alpha_{s}(u)=1+\frac{i}{u}d_{s}+\mathcal{O}(u^{-2}),$ (S-10)
for $s=1,\dots,n$. The $d_{s}$ are the Dynkin labels of the $\mathfrak{g}$-representation corresponding to the pseudo-vacuum. We already showed a compatibility condition for the Hilbert space, namely $\alpha_{s}(u)\alpha_{\rho(s)}(-u)=1$. For chiral nodes it does not give any restriction for the Dynkin label $d_{s}$ but it does for the achiral ones. For the achiral $\mathfrak{gl}_{N}$ MPS the Hilbert space has to satisfy $d_{N-s}=d_{s}$ where $s=1,\dots,N$ and for the achiral $\mathfrak{o}_{2n}$ MPS the Hilbert space has to satisfy $d_{n-1}=d_{n}$.
In the following, we go through the possible reflection algebras and the corresponding extra selection rules of the on-shell overlaps.
$Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$ (chiral)
For the $Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$ there is no extra selection rule, therefore the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing.
$Y(\mathfrak{gl}_{2k},\mathfrak{sp}_{2k})$ (chiral)
For the $Y(\mathfrak{gl}_{2k},\mathfrak{sp}_{2k})$ the overlap is non-vanishing only when the numbers $r_{2a}$ are even for $a=1,\dots,k-1$, and
$r_{2a-1}=\frac{d_{2a-1}+r_{2a-2}+r_{2a}}{2},$ (S-11)
for $a=1,\dots,k$. We can see that the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing only when $d_{2a-1}=0$ for $a=1,\dots,k$.
$Y(\mathfrak{gl}_{N},\mathfrak{gl}_{M}\oplus\mathfrak{gl}_{N-M})$ (achiral)
For the $Y(\mathfrak{gl}_{2n},\mathfrak{gl}_{n}\oplus\mathfrak{gl}_{n})$ or $Y(\mathfrak{gl}_{2n+1},\mathfrak{gl}_{n}\oplus\mathfrak{gl}_{n+1})$ the overlaps do not have extra selection rules. For $M<\left\lfloor\frac{N}{2}\right\rfloor$ the overlap is non-vanishing only when
$r_{j}=r_{M}+\sum_{k=M+1}^{j}\Lambda_{k},$ (S-12)
for $j=M+1,\dots,n_{+}$, where we introduced the weights
$\Lambda_{j}=\begin{cases}\sum_{k=j}^{n-1}d_{k}+\frac{1}{2}d_{n},&N=2n,\\ \sum_{k=j}^{n}d_{k},&N=2n+1,\end{cases}$ (S-13)
for $j=M+1,\dots,n_{+}$. We can see that the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing only when $d_{s}=0$ for $s=M+1,\dots,N-M$.
$Y(\mathfrak{o}_{2n+1},\mathfrak{o}_{M}\oplus\mathfrak{o}_{2n+1-M})$ (chiral)
[…]
$r_{j}=r_{M}+\sum_{k=M+1}^{j}\Lambda_{k},$ (S-14)
for $j=M+1,\dots,n$, where we introduced the weights
$\Lambda_{j}=\sum_{k=j}^{n}d_{k},$ (S-15)
for $j=M+1,\dots,n$. We can see that the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing only when $d_{s}=0$ for $s=M+1,\dots,n$.
$Y(\mathfrak{sp}_{2n},\mathfrak{gl}_{n})$ (chiral)
For the $Y(\mathfrak{sp}_{2n},\mathfrak{gl}_{n})$ there is no extra selection rule, therefore the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing.
$Y(\mathfrak{sp}_{2n},\mathfrak{sp}_{2m}\oplus\mathfrak{sp}_{2n-2m})$ (chiral)
[…]
$\begin{split}r_{2a-1}&=\frac{r_{2a-2}+r_{2a}+d_{2a-1}}{2},\quad\text{for }a=1,\dots,m,\\ r_{j}&=r_{2m}+\sum_{k=2m+1}^{j}\Lambda_{k},\quad j=2m+1,\dots,n-1,\\ r_{n}&=\frac{1}{2}r_{2m}+\frac{1}{2}\sum_{k=2m+1}^{n}\Lambda_{k},\end{split}$ (S-16)
where we introduced the weights
$\Lambda_{j}=\sum_{k=j}^{n-1}d_{k}+\frac{1}{2}d_{n},$ (S-17)
for $j=M+1,\dots,n$. We can see that the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing only when $d_{2a-1}=d_{s}=0$ for $a=1,\dots,m$ and $s=2m+1,\dots,n$.
$Y(\mathfrak{o}_{2n},\mathfrak{o}_{M}\oplus\mathfrak{o}_{2n-M})$
[…]
$\begin{split}r_{j}&=r_{M}+\sum_{k=M+1}^{j}\Lambda_{k},\quad j=M+1,\dots,n-2,\\ r_{n-1}&=\frac{1}{2}r_{M}+\frac{1}{2}\sum_{k=M+1}^{n-1}\Lambda_{k}-\frac{1}{2}\Lambda_{n},\\ r_{n}&=\frac{1}{2}r_{M}+\frac{1}{2}\sum_{k=M+1}^{n-1}\Lambda_{k}+\frac{1}{2}\Lambda_{n},\end{split}$ (S-18)
where we introduced the weights
$\begin{split}\Lambda_{j}&=\sum_{k=j}^{n-2}d_{k}+\frac{1}{2}d_{n-1}+\frac{1}{2}d_{n},\quad j=M+1,\dots,n-1,\\ \Lambda_{n}&=\frac{1}{2}d_{n}-\frac{1}{2}d_{n-1}.\end{split}$ (S-19)
We can see that the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing only when $d_{s}=0$ for $s=M+1,\dots,n$.
$Y(\mathfrak{o}_{2n},\mathfrak{gl}_{n})$
If $n$ is even (chiral) then
$r_{2a-1}=\frac{r_{2a-2}+r_{2a}+d_{2a-1}}{2},$ (S-20)
for $a=1,\dots,\frac{n}{2}.$ When $n$ is odd (achiral) the selection rules are
$\begin{split}r_{2a-1}&=\frac{r_{2a-2}+r_{2a}+d_{2a-1}}{2},\quad\text{for }a=1,\dots,\frac{n-3}{2},\\ r_{n-2}&=\frac{r_{n-3}+2r_{n-1}+d_{n-2}}{2}.\end{split}$ (S-21)
We can see that the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing only when $d_{2a-1}=0$ for $a=1,\dots,\frac{n}{2}$ when $n$ is even and for $a=1,\dots,\frac{n-1}{2}$ when $n$ is odd.
Comments on the $\beta$-parameters of the MPS overlap function
We just saw that for certain Hilbert spaces, the extra selection rules lead to zero pseudo-vacuum overlap, i.e., $\mathbf{B}=\langle\Psi|0\rangle=0$. Our general overlap formula (S-1) is still valid, but for these Hilbert spaces the $\beta$-parameters are not eigenvalues of the operator $\mathbf{B}$. In these cases, the closed formula for the $\beta$-parameters is not known (yet). Since these are simple numbers (do not depending on the Bethe roots), their determination is not too complicated in the concrete cases. They can be determined by a simple numerical analysis.
To summarize, for Hilbert spaces where the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-zero, the $\beta$-parameters are eigenvalues of $\mathbf{B}$. In the cases where the operator B is zero the determination of $\beta_{k}$-s is a separate problem.
S-III $F$-operators for $\mathfrak{g}=\mathfrak{gl}_{N}$
In the $\mathfrak{g}=\mathfrak{gl}_{N}$ case there exist chiral or achiral pair structures.
For chiral pair structure we introduce $N$ commuting operators $\mathbf{G}^{(k)}(u)$, for $k=1,\dots,N$. We also introduce $F$-operators
$\mathbf{F}^{(k)}(u)=\left[\mathbf{G}^{(k)}(u)\right]^{-1}\mathbf{G}^{(k+1)}(u),$ (S-22)
for $k=1,\dots,N-1$. They can be diagonalized simultaneously
$\mathbf{F}^{(k)}(u)=\mathbf{A}\mathrm{diag}(\mathcal{F}_{1}^{(k)}(u),\dots,\mathcal{F}_{d_{B}}^{(k)}(u))\mathbf{A}^{-1},$ (S-23)
where $\mathbf{A}$ is an invertible matrix of the boundary space. The final $\tilde{\mathcal{F}}$ functions are defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\mathcal{F}_{j}^{(k)}(iu-k/2)\sqrt{\frac{u^{2}}{u^{2}+1/4}}.$ (S-24)
For achiral pair structure we introduce $n_{+}+1=\left\lfloor\frac{N}{2}\right\rfloor+1$ commuting operators $\mathbf{G}^{(k)}(u)$, for $k=1,\dots,n_{+}+1$. The $F$-operators and their spectrum are defined in the same way as before (S-22), (S-23) for $k=1,\dots,n_{+}$. The final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu+\frac{N-2k}{4}),&k<\frac{N-1}{2},\\ \mathcal{F}_{j}^{(\frac{N-1}{2})}(iu+\frac{1}{4})\sqrt{\frac{u-i/4}{u+i/4}},&k=\frac{N-1}{2},\\ \mathcal{F}_{j}^{(\frac{N}{2})}(iu)\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=\frac{N}{2}.\end{cases}$ (S-25)
The pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ also commutes with the $F$-operators and its spectrum gives the $\beta_{j}$-s of the overlap formula
$\mathbf{B}=\mathbf{A}\mathrm{diag}(\beta_{1},\dots,\beta_{d_{B}})\mathbf{A}^{-1}.$ (S-26)
S-III.1 $Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$ (chiral)
For the twisted Yangian $Y^{+}(N)\equiv Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$ we define a series of nested K-matrices with the following recursion
$\mathbf{K}_{a,b}^{(k+1)}(u)=\mathbf{K}_{a,b}^{(k)}(u)-\mathbf{K}_{a,k}^{(k)}(u)\left[\mathbf{K}_{k,k}^{(k)}(u)\right]^{-1}\mathbf{K}_{k,b}^{(k)}(u),$ (S-27)
where $a,b=k+1,\dots,N$. The starting condition is $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We define the commuting operators as
$\mathbf{G}^{(k)}(u)=\mathbf{K}_{k,k}^{(k)}(u),$ (S-28)
for $k=1,\dots,N$.
S-III.2 $Y(\mathfrak{gl}_{2k},\mathfrak{sp}_{2k})$ (chiral)
S-III.2.1 $k=1$
We start with the $k=1$ case, i.e. $Y^{-}(2)\equiv Y(\mathfrak{gl}_{2},\mathfrak{sp}_{2})$. At first we have to find the factorization form
$\mathbf{K}_{a,b}(u)=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u)\epsilon_{c,d}\mathbf{L}_{d,b}(-u).$ (S-29)
We can define the deformed K-matrix
$\tilde{\mathbf{K}}_{a,b}(u)=\mathbf{K}_{a,b}(u)+(u+1/2)\mathbf{L}_{1,a}(u)\mathbf{L}_{1,b}(-u).$ (S-30)
Now we can define the G-operators
$\begin{split}\mathbf{G}^{(1)}(u)&=\tilde{\mathbf{K}}_{1,1}(u),\\ \mathbf{G}^{(2)}(u)&=\tilde{\mathbf{K}}_{2,2}(u)-\tilde{\mathbf{K}}_{2,1}(u)\left[\tilde{\mathbf{K}}_{1,1}(u)\right]^{-1}\tilde{\mathbf{K}}_{1,2}(u).\end{split}$ (S-31)
S-III.2.2 General $k$
For general $k$, i.e. $Y^{-}(2k)\equiv Y(\mathfrak{gl}_{2k},\mathfrak{sp}_{2k})$, the nesting is defined as
$\mathbf{K}_{a,b}^{(s+1)}(u)=\mathbf{K}_{a,b}^{(s)}(u)-\sum_{\alpha,\beta=2s-1}^{2s}\mathbf{K}_{a,\alpha}^{(s)}(u)\widehat{\mathbf{K}}_{\alpha,\beta}^{(s)}(u)\mathbf{K}_{\beta,b}^{(s)}(u),$ (S-32)
for $a,b=2s+1,\dots,2k$. We defined the inverse matrix $\widehat{\mathbf{K}}_{\alpha,\beta}^{(s)}(u)$ for $\alpha,\beta=2s-1,2s$ as
$\sum_{\beta=2s-1}^{2s}\widehat{\mathbf{K}}_{\alpha,\beta}^{(s)}(u)\mathbf{K}_{\beta,\gamma}^{(s)}(u)=\delta_{\alpha,\gamma}.$ (S-33)
We can select $k$ commuting $Y^{-}(2)$ subalgebras
$\mathbf{k}^{(s)}(u)=\left(\begin{array}[]{cc}\mathbf{K}_{2s-1,2s-1}^{(s)}(u)&\mathbf{K}_{2s-1,2s}^{(s)}(u)\\ \mathbf{K}_{2s,2s-1}^{(s)}(u)&\mathbf{K}_{2s,2s}^{(s)}(u)\end{array}\right),$ (S-34)
where $s=1,\dots,k$. We can repeat the $Y^{-}(2)$ method separately. We have to find the factorization form
$\mathbf{k}_{a,b}^{(s)}(u)=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u+s-1)\epsilon_{c,d}\mathbf{L}_{d,b}(-u-s+1),$ (S-35)
and the deformed K-matrix
$\tilde{\mathbf{k}}_{a,b}^{(s)}(u)=\mathbf{k}_{a,b}^{(s)}(u)+(u+s-1/2)\mathbf{L}_{1,a}(u+s-1)\mathbf{L}_{1,b}(-u-s+1).$ (S-36)
Now we can define the G-operators
$\begin{split}\mathbf{G}^{(2s-1)}(u)&=\tilde{\mathbf{k}}_{1,1}^{(s)}(u),\\ \mathbf{G}^{(2s)}(u)&=\tilde{\mathbf{k}}_{2,2}^{(s)}(u)-\tilde{\mathbf{k}}_{2,1}^{(s)}(u)\left[\tilde{\mathbf{k}}_{1,1}^{(s)}(u)\right]^{-1}\tilde{\mathbf{k}}_{1,2}^{(s)}(u),\end{split}$ (S-37)
for $s=1,\dots,k$.
S-III.3 $Y(\mathfrak{gl}_{N},\mathfrak{gl}_{M}\oplus\mathfrak{gl}_{N-M})$ (achiral)
In this subsection, we use a symmetric convention for the indices $a,b$ of the the K-matrix $\mathbf{K}_{a,b}$. For odd $N=2n+1$ the indices $a,b$ run through the set $\left\{-n,...,-1,0,1,...,n\right\}$, i.e. the K-matrix has the following components
$\mathbf{K}=\left(\begin{array}[]{ccccccc}\mathbf{K}_{-n,-n}&\dots&\mathbf{K}_{-n,-1}&\mathbf{K}_{-n,0}&\mathbf{K}_{-n,1}&\dots&\mathbf{K}_{-n,n}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\iddots&\vdots\\ \mathbf{K}_{-1,-n}&\dots&\mathbf{K}_{-1,-1}&\mathbf{K}_{-1,0}&\mathbf{K}_{-1,1}&\dots&\mathbf{K}_{-1,n}\\ \mathbf{K}_{0,-n}&\dots&\mathbf{K}_{0,-1}&\mathbf{K}_{0,0}&\mathbf{K}_{0,1}&\dots&\mathbf{K}_{0,n}\\ \mathbf{K}_{1,-n}&\dots&\mathbf{K}_{1,-1}&\mathbf{K}_{1,0}&\mathbf{K}_{1,1}&\dots&\mathbf{K}_{1,n}\\ \vdots&\iddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{K}_{n,-n}&\dots&\mathbf{K}_{n,-1}&\mathbf{K}_{n,0}&\mathbf{K}_{n,1}&\dots&\mathbf{K}_{n,n}\end{array}\right).$ (S-38)
For even $N=2n$ the indices $a,b$ run through the set $\left\{-n,...,-1,1,...,n\right\}$, i.e. the K-matrix has the following components
$\mathbf{K}=\left(\begin{array}[]{cccccc}\mathbf{K}_{-n,-n}&\dots&\mathbf{K}_{-n,-1}&\mathbf{K}_{-n,1}&\dots&\mathbf{K}_{-n,n}\\ \vdots&\ddots&\vdots&\vdots&\iddots&\vdots\\ \mathbf{K}_{-1,-n}&\dots&\mathbf{K}_{-1,-1}&\mathbf{K}_{-1,1}&\dots&\mathbf{K}_{-1,n}\\ \mathbf{K}_{1,-n}&\dots&\mathbf{K}_{1,-1}&\mathbf{K}_{1,1}&\dots&\mathbf{K}_{1,n}\\ \vdots&\iddots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{K}_{n,-n}&\dots&\mathbf{K}_{n,-1}&\mathbf{K}_{n,1}&\dots&\mathbf{K}_{n,n}\end{array}\right).$ (S-39)
S-III.3.1 $Y(\mathfrak{gl}_{2n},\mathfrak{gl}_{n}\oplus\mathfrak{gl}_{n})$ or $Y(\mathfrak{gl}_{2n+1},\mathfrak{gl}_{n}\oplus\mathfrak{gl}_{n+1})$
When $M=\left\lfloor\frac{N}{2}\right\rfloor$ we have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,n$ with the recursion
$\mathbf{K}_{a,b}^{(s+1)}(u)=\mathbf{K}_{a,b}^{(s)}(u)-\mathbf{K}_{a,-n+s-1}^{(s)}(u)\left[\mathbf{K}_{n-s+1,-n+s-1}^{(s)}(u)\right]^{-1}\mathbf{K}_{n-s+1,b}^{(s)}(u),$ (S-40)
where $a,b=-n+s,-n+s+1,\dots,n-s-1,n-s$, with the initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,n$ with
$\mathbf{G}^{(s)}(u)=\mathbf{K}_{n-s+1,-n+s-1}^{(s)}(u),$ (S-41)
and
$\mathbf{G}^{(n+1)}(u)=\begin{cases}\mathbf{K}_{0,0}^{(n)}(u)-\mathbf{K}_{0,-1}^{(n)}(u)\left[\mathbf{K}_{1,-1}^{(n)}(u)\right]^{-1}\mathbf{K}_{1,0}^{(n)}(u),&\text{for }N=2n+1,\\ \mathbf{K}_{-1,1}^{(n)}(u)-\mathbf{K}_{-1,-1}^{(n)}(u)\left[\mathbf{K}_{1,-1}^{(n)}(u)\right]^{-1}\mathbf{K}_{1,1}^{(n)}(u),&\text{for }N=2n.\end{cases}$ (S-42)
S-III.3.2 $M=0$
For $M=0$. at first we have to find the factorization form
$\mathbf{K}(u)=\mathbf{L}^{-1}(u)\mathbf{L}(-u).$ (S-43)
We can define the deformed K-matrix
$\tilde{\mathbf{K}}_{a,b}(u)=\mathbf{K}_{a,b}(u)+u\sum_{c=1}^{n}\left[\mathbf{L}^{-1}(u)\right]_{a,c}\mathbf{L}_{-c,b}(-u).$ (S-44)
Now we can use the nesting (S-40) with the initial condition $\mathbf{K}^{(1)}(u)\equiv\tilde{\mathbf{K}}(u)$ to obtain the nested K-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,n$. We define the $G$-operators with (S-41), (S-42).
S-III.3.3 General $M$
For general $M$ we combine the methods of sections S-III.3.1 and S-III.3.2. We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,M+1$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,M$ with (S-41).
The K-matrix $\mathbf{K}^{(M+1)}$ is a $Y(\mathfrak{gl}_{N-2M},\mathfrak{gl}_{N-2M})$ representation therefore we continue with the method of S-III.3.2. We have to find the factorization form
$\mathbf{K}^{(M+1)}(u)=\mathbf{L}^{-1}(u)\mathbf{L}(-u),$ (S-45)
and deform the K-matrix
$\tilde{\mathbf{K}}_{a,b}^{(M+1)}(u)=\mathbf{K}_{a,b}^{(M+1)}(u)+u\sum_{c=1}^{n-M}\left[\mathbf{L}^{-1}(u)\right]_{a,c}\mathbf{L}_{-c,b}(-u).$ (S-46)
We can continue the series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=M+2,\dots,n$ with the recursion (S-40) and initial condition $\mathbf{K}^{(M+1)}(u)\equiv\tilde{\mathbf{K}}^{(M+1)}$. We define the remaining $G$-operators as $\mathbf{G}^{(M+1)}(u)=\tilde{\mathbf{K}}_{n-M,-n+M}^{(M+1)}(u)$ and with the equation (S-41) for $s=M+2,M+3,\dots,n$ and the $\mathbf{G}^{(n+1)}(u)$ with (S-42).
S-IV $F$-operators for $\mathfrak{g}=\mathfrak{o}_{2n}$
For $\mathfrak{g}=\mathfrak{o}_{2n}$ there exists chiral or achiral pair structure. For the chiral case, we introduce $n+1$ commuting operators $\mathbf{G}^{(k)}(u)$, for $k=1,\dots,n+1$. We also introduce $F$-operators
$\mathbf{F}^{(k)}(u)=\begin{cases}\left[\mathbf{G}^{(k)}(u)\right]^{-1}\mathbf{G}^{(k+1)}(u),&k=1,\dots,n-1,\\ \left[\mathbf{G}^{(n-1)}(u)\right]^{-1}\mathbf{G}^{(n+1)}(u),&k=n,\end{cases}$ (S-47)
and the final $\tilde{\mathcal{F}}$ functions as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu+\frac{n-k-1}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-2,\\ \mathcal{F}_{j}^{(k)}(iu)\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=n-1,n.\end{cases}$ (S-48)
For achiral pair structure we introduce $n$ commuting operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n$. The $F$-operators are defined as (S-22) for $k=1,\dots,n-1$. The final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu+\frac{n-k-1}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-2,\\ \mathcal{F}_{j}^{(n-1)}(iu)\frac{u}{u+i/2},&k=n-1.\end{cases}$ (S-49)
In this section, we use the convention of section S-III.3 for the indices $a,b$ of the K-matrix $\mathbf{K}_{a,b}$, i.e., the indices $a,b$ run through the set $\left\{-n,...,-1,1,...,n\right\}$, and the K-matrix has the components (S-39).
S-IV.1 $Y(\mathfrak{o}_{2n},\mathfrak{gl}_{n})$
S-IV.1.1 $Y(\mathfrak{o}_{4k},\mathfrak{gl}_{2k})$ (chiral)
We define the nested $K$-matrices $\mathbf{K}^{(s)}$ for $s=2,\dots,k$ ($\mathbf{K}^{(1)}\equiv\mathbf{K}$) with the recursion
$\mathbf{K}_{a,b}^{(s+1)}(u)=\mathbf{K}_{a,b}^{(s)}(u)-\sum_{\alpha,\beta=n-2s+1}^{n-2s+2}\mathbf{K}_{a,-\alpha}^{(s)}(u)\widehat{\mathbf{K}}_{-\alpha,\beta}^{(s)}(u)\mathbf{K}_{\beta,b}^{(s)}(u),$ (S-50)
where $a,b=-n+2s,-n+2s+1,\dots,n-2s-1,n-2s$. We defined the inverse matrix $\widehat{\mathbf{K}}_{-\alpha,\beta}^{(s)}(u)$ for $\alpha,\beta=n-2s+1,n-2s+2$ as
$\sum_{\beta=n-2s+1}^{n-2s+2}\widehat{\mathbf{K}}_{-\alpha,\beta}^{(s)}(u)\mathbf{K}_{\beta,-\gamma}^{(s)}(u)=\delta_{\alpha,\gamma},$ (S-51)
where $s=1,\dots,k$. We can select commuting $Y^{-}(2)$ subalgebras
$\mathbf{k}^{(s)}(u)=\left(\begin{array}[]{cc}\mathbf{K}_{n-2s+2,-n+2s-2}^{(s)}(u)&\mathbf{K}_{n-2s+2,-n+2s-1}^{(s)}(u)\\ \mathbf{K}_{n-2s+1,-n+2s-2}^{(s)}(u)&\mathbf{K}_{n-2s+1,-n+2s-1}^{(s)}(u)\end{array}\right),$ (S-52)
for $s=1,\dots,k$. We can use the $Y^{-}(2)$ method. We have to find the deformed $K$-matrices
$\begin{split}\mathbf{k}_{a,b}^{(s)}(u)&=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u-\kappa_{N-4s+4})\epsilon_{c,d}\mathbf{L}_{d,b}(-u),\\ \tilde{\mathbf{k}}_{a,b}^{(s)}(u)&=\mathbf{k}_{a,b}^{(s)}(u)+(u+1/2-\kappa_{N-4s+4}/2)\mathbf{L}_{1,a}(u-\kappa_{N-4s+4})\mathbf{L}_{1,b}(-u),\end{split}$ (S-53)
for $s=1,\dots,k$. The G-operators $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(n)}$ are defined with (S-37) for $s=1,\dots,k$. They define the F-operators in the usual way
$\mathbf{F}^{(s)}(u)=\left[\mathbf{G}^{(s)}(u)\right]^{-1}\mathbf{G}^{(s+1)}(u),$ (S-54)
for $s=1,\dots,n-1$. The remaining $F$-operator is
$\mathbf{F}^{(n)}(u)=\left[\tilde{\mathbf{K}}_{2,-1}^{(k)}(u)\right]^{-1}\left[\tilde{\mathbf{K}}_{-1,2}^{(k)}(u)-\tilde{\mathbf{K}}_{-1,-1}^{(k)}(u)\left[\tilde{\mathbf{K}}_{2,-1}^{(k)}(u)\right]^{-1}\tilde{\mathbf{K}}_{2,2}^{(k)}(u)\right].$ (S-55)
S-IV.1.2 $Y(\mathfrak{o}_{4k+2},\mathfrak{gl}_{2k+1})$ (achiral)
We define the nested $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,k$ with the recursion (S-50) and initial condition $\mathbf{K}^{(1)}\equiv\mathbf{K}$. We also define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53). The G-operators $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(n-1)}$ are defined with (S-37) for $s=1,\dots,k$. The remaining $G$-operator is
$\mathbf{G}^{(n)}(u)=\mathbf{K}_{1,1}^{(k)}(u)-\sum_{\alpha,\beta=2}^{3}\mathbf{K}_{1,-\alpha}^{(k)}(u)\widehat{\mathbf{K}}_{-\alpha,\beta}^{(k)}(u)\mathbf{K}_{\beta,1}^{(k)}(u).$ (S-56)
S-IV.2 $Y(\mathfrak{o}_{2n},\mathfrak{o}_{M}\oplus\mathfrak{o}_{2n-M})$
S-IV.2.1 $Y(\mathfrak{o}_{2n},\mathfrak{o}_{n}\oplus\mathfrak{o}_{n})$ (chiral)
We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,n-1$ with the recursion (S-40). We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,n-1$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(n)}(u),\mathbf{G}^{(n+1)}(u)$ as
$\begin{split}\mathbf{G}^{(n)}(u)&=\mathbf{K}_{1,-1}^{(n-1)}(u)-\mathbf{K}_{1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,-1}^{(n-1)}(u),\\ \mathbf{G}^{(n+1)}(u)&=\mathbf{K}_{-1,1}^{(n-1)}(u)-\mathbf{K}_{-1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,1}^{(n-1)}(u).\end{split}$ (S-57)
S-IV.2.2 $Y(\mathfrak{o}_{2n},\mathfrak{o}_{n-1}\oplus\mathfrak{o}_{n+1})$ (achiral)
We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,n-1$ with the recursion (S-40). We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,n-1$ with (S-41). We also need to define the remaining $G$-operator $\mathbf{G}^{(n)}(u)$ as
$\mathbf{G}^{(n)}(u)=\mathbf{K}_{1,1}^{(n-1)}(u)-\mathbf{K}_{1,-2}^{(n-1)}(u)\left[\mathbf{K}_{2,-2}^{(n-1)}(u)\right]^{-1}\mathbf{K}_{2,1}^{(n-1)}(u).$ (S-58)
S-IV.2.3 General $M$
We continue with general $n$ and $M$. We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,M+1$ with the recursion (S-40). We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,M$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$ for the chiral and $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n)}(u)$ for achiral cases. When $n-M$ is even the pair structure is chiral and the pair structure is odd when $n-M$ is odd.
If $n-M$ is even, i.e. $n-M=2k$ ($k\in\mathbb{Z}_{+}$) we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ as
$\begin{split}\mathbf{K}_{a,b}^{(M+1)}(u)&=\sum_{c=-\tilde{n}}^{\tilde{n}}\mathbf{L}_{-c,-a}(u-\kappa_{N-2M})\mathbf{L}_{c,b}(-u),\\ \tilde{\mathbf{K}}_{a,b}^{(1)}(u)&=\mathbf{K}_{a,b}^{(M+1)}(u)+\\ &+u\sum_{l=1}^{k}\left(\mathbf{L}_{-(\tilde{n}-2l+1),-a}(u-\kappa_{N-2M})\mathbf{L}_{-\tilde{n}+2l-2,b}(-u)-\mathbf{L}_{-(\tilde{n}-2l+2),-a}(u-\kappa_{N-2M})\mathbf{L}_{-\tilde{n}+2l-1,b}(-u)\right),\end{split}$ (S-59)
where $\tilde{n}=2k=n-M$. We continue with the series of the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k$ as (S-50) and define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n)}$ are
$\begin{split}\mathbf{G}^{(M+2s-1)}(u)&=\tilde{\mathbf{k}}_{1,1}^{(s)}(u),\\ \mathbf{G}^{(M+2s)}(u)&=\tilde{\mathbf{k}}_{2,2}^{(s)}(u)-\tilde{\mathbf{k}}_{2,1}^{(s)}(u)\left[\tilde{\mathbf{k}}_{1,1}^{(s)}(u)\right]^{-1}\tilde{\mathbf{k}}_{1,2}^{(s)}(u),\end{split}$ (S-60)
for $s=1,\dots,k$. The F-operators are defined as (S-54) and (S-55).
If $n-M$ is odd, i.e. $n-M=2k+1$ ($k\in\mathbb{Z}_{+}$) we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ as (S-59) where $\tilde{n}=2k+1=n-M$. We continue with the series of the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k$ as (S-50). We define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n-1)}$ are defined with (S-60) for $s=1,\dots,k$. The remaining $G$-operator $\mathbf{G}^{(n)}$ is defined as (S-56).
S-V $F$-operators for $\mathfrak{g}=\mathfrak{o}_{2n+1}$
In the $\mathfrak{g}=\mathfrak{o}_{2n+1}$ case there exists only chiral pair structure. We introduce $n+1$ $G$-operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n+1$ and we introduce the $F$-operators as (S-22). The final $\tilde{\mathcal{F}}$ functions are defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu-\frac{1}{4}+\frac{n-k}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-1,\\ \mathcal{F}_{j}^{(n)}(iu-\frac{1}{4})\frac{u+i/4}{u+i/2}\sqrt{\frac{u^{2}}{u^{2}+1/16}},&k=n.\end{cases}$ (S-61)
In this section, we use the convention of section S-III.3 for the indices $a,b$ of the K-matrix $\mathbf{K}_{a,b}$, i.e., the indices $a,b$ run through the set $\left\{-n,...,-1,0,1,...,n\right\}$, and the K-matrix has the components (S-38).
S-V.1 $Y(\mathfrak{o}_{2n+1},\mathfrak{o}_{n}\oplus\mathfrak{o}_{n+1})$
For $Y(\mathfrak{o}_{2n+1},\mathfrak{o}_{n}\oplus\mathfrak{o}_{n+1})$ we have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,n$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,n$ with (S-41) and $\mathbf{G}^{(n+1)}$ with (S-42).
S-V.2 $Y(\mathfrak{o}_{2n+1},\mathfrak{o}_{M}\oplus\mathfrak{o}_{2n-M+1})$
For general $M$ we have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,M+1$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,M$ with (S-41). We also need to define the remaining $G$-operators $\mathbf{G}^{(M+1)}(u),\mathbf{G}^{(M+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$. Their definitions depend on whether $n-M$ is even or odd.
If $n-M$ is odd, i.e. $n-M=2k+1$ ($k\in\mathbb{N}$) we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ with (S-59) where $\tilde{n}=2k+1=n-M$. We define the nested $K$-matrices $\tilde{\mathbf{K}}_{a,b}^{(s)}(u)$ for $s=2,\dots,k+1$ with (S-50), the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53) (we replace $n\to\tilde{n}$ , $\mathbf{K}^{(s)}\to\tilde{\mathbf{K}}^{(s)}$ and $\kappa_{N-4s+4}\to\kappa_{N-2M-4s+4}$). We can also define the G-operators $\mathbf{G}^{(r)}$ for $r=M+1,\dots,n-1$ as (S-60) for $s=1,\dots,k$. We also have to deform the K-matrix $\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)$ as
$\begin{split}\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)&=\sum_{c=-1}^{1}\mathbf{L}_{-c,-a}(u-\kappa_{3})\mathbf{L}_{c,b}(-u),\\ \check{\mathbf{K}}_{a,b}^{(k+1)}(u)&=\tilde{\mathbf{K}}_{a,b}^{(k+1)}(u)-\frac{1}{2}u(u+\frac{1}{4})\mathbf{L}_{-1,-a}(u-\kappa_{3})\mathbf{L}_{-1,b}(-u)\\ &+u\left(\mathbf{L}_{0,-a}(u-\kappa_{3})\mathbf{L}_{-1,b}(-u)-\mathbf{L}_{-1,-a}(u-\kappa_{3})\mathbf{L}_{0,b}(-u)\right).\end{split}$ (S-62)
and the remaining $G$-operators $\mathbf{G}^{(n)}$, $\mathbf{G}^{(n+1)}$ are
$\begin{split}\mathbf{G}^{(n)}(u)&=\check{\mathbf{K}}_{1,-1}^{(k+1)}(u),\\ \mathbf{G}^{(n+1)}(u)&=\check{\mathbf{K}}_{0,0}^{(k+1)}(u)-\check{\mathbf{K}}_{0,-1}^{(k+1)}(u)\left[\check{\mathbf{K}}_{1,-1}^{(k+1)}(u)\right]^{-1}\check{\mathbf{K}}_{1,0}^{(k+1)}(u).\end{split}$ (S-63)
If $n-M$ is even, i.e. $n-M=2k$ ($k\in\mathbb{Z}_{+}$), we have to deform the K-matrix $\mathbf{K}^{(M+1)}$ with (S-59) where we replace $n\to\tilde{n}=2k=n-M$ , $\mathbf{K}\to\mathbf{K}^{(M+1)}$ and $\kappa_{N}\to\kappa_{N-2M}$. We also define the nested $K$-matrices $\tilde{\mathbf{K}}^{(s)}$ for $s=1,\dots,k$ with the recursion (S-50) and the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,k$ with (S-53). The G-operators $\mathbf{G}^{(M+1)},\dots,\mathbf{G}^{(n)}$ are defined with (S-60) for $s=1,\dots,k$. The remaining $G$-operator $\mathbf{G}^{(n+1)}$ is defined as
$\mathbf{G}^{(n+1)}(u)=\mathbf{K}_{0,0}^{(k)}(u)-\sum_{\alpha,\beta=1}^{2}\mathbf{K}_{0,-\alpha}^{(k)}(u)\widehat{\mathbf{K}}_{-\alpha,\beta}^{(k)}(u)\mathbf{K}_{\beta,0}^{(k)}(u).$ (S-64)
S-VI $F$-operators for $\mathfrak{g}=\mathfrak{sp}_{2n}$
In the $\mathfrak{g}=\mathfrak{sp}_{2n}$ case there exists only chiral pair structure. We introduce $n+1$ G-operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n+1$ and we introduce the $F$-operators as (S-22) and the final $\tilde{\mathcal{F}}$ functions are defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu+\frac{n-k}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-1,\\ \mathcal{F}_{j}^{(n)}(iu)\sqrt{\frac{u^{2}}{u^{2}+1}},&k=n.\end{cases}$ (S-65)
In this section, we use the convention of section S-III.3 for the indices $a,b$ of the K-matrix $\mathbf{K}_{a,b}$, i.e., the indices $a,b$ run through the set $\left\{-n,...,-1,1,...,n\right\}$, and the K-matrix has the components (S-39).
S-VI.1 $Y(\mathfrak{sp}_{2n},\mathfrak{gl}_{n})$
We have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,n$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,n$ with (S-41) and $\mathbf{G}^{(n+1)}$ with (S-42).
S-VI.2 $Y(\mathfrak{sp}_{2n},\mathfrak{sp}_{2m}\oplus\mathfrak{sp}_{2n-2m})$
We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,m+1$ with the recursion (S-50) and initial condition $\mathbf{K}^{(1)}\equiv\mathbf{K}$. We also define the $Y^{-}(2)$ $K$-matrices $\mathbf{k}^{(s)}$ and the deformed ones $\tilde{\mathbf{k}}^{(s)}$ for $s=1,\dots,m$ with (S-53). We can define the G-operators $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(2m)}$ with (S-37) for $s=1,\dots,m$. For $n>2m$ we also need to define the remaining $G$-operators $\mathbf{G}^{(2m+1)}(u),\mathbf{G}^{(2m+2)}(u),\dots,\mathbf{G}^{(n+1)}(u)$.
For $k=n-2m>0$, we have to deform the K-matrix $\mathbf{K}^{(2m+1)}$ as
$\begin{split}\mathbf{K}_{a,b}^{(2m+1)}(u)&=\sum_{c=-k}^{k}\theta_{a}\theta_{c}\mathbf{L}_{-c,-a}(u-\kappa_{2k})\mathbf{L}_{c,b}(-u),\\ \tilde{\mathbf{K}}_{a,b}^{(1)}(u)&=\mathbf{K}_{a,b}^{(2m+1)}(u)+\sum_{c=1}^{k}\theta_{a}\theta_{c}\mathbf{L}_{-c,-a}(u-\kappa_{2k})\mathbf{L}_{-c,b}(-u),\end{split}$ (S-66)
where
$\theta_{i}=\begin{cases}+1,&i>0,\\ -1,&i<0.\end{cases}$ (S-67)
We also define the nested $K$-matrices $\tilde{\mathbf{K}}^{(s)}$ for $s=2,\dots,k$ with the recursion (S-40). The G-operators $\mathbf{G}^{(2m+1)},\dots,\mathbf{G}^{(n)}$ are defined as
$\mathbf{G}^{(2m+s)}(u)=\tilde{\mathbf{K}}_{k-s+1,-k+s-1}^{(s)}(u),$ (S-68)
for $s=1,\dots,k$ and
$\mathbf{G}^{(n+1)}(u)=\tilde{\mathbf{K}}_{-1,1}^{(k)}(u)-\tilde{\mathbf{K}}_{-1,-1}^{(k)}(u)\left[\tilde{\mathbf{K}}_{1,-1}^{(k)}(u)\right]^{-1}\tilde{\mathbf{K}}_{1,1}^{(k)}(u).$ (S-69)
S-VII Examples
In this section, we demonstrate the application of the overlap formulas on concrete examples.
S-VII.1 $SO(6)$ spin chain
Let us consider the $SO(6)$ spin chain which is given by the Hamiltonian
$H=\sum_{j=1}^{J}\mathbf{P}_{j,j+1}-\frac{1}{2}\mathbf{Q}_{j,j+1},$ (S-70)
where
$\mathbf{P}=\sum_{i,j=1}^{6}e_{i,j}\otimes e_{j,i},\quad\mathbf{Q}=\sum_{i,j=1}^{6}e_{i,j}\otimes e_{7-i,7-j}.$ (S-71)
In the AdS/CFT context we often identify the states as the scalar fields of the $\mathcal{N}=4$ SYM as
$|1\rangle\equiv Z,\quad|2\rangle\equiv Y,\quad|3\rangle\equiv X,\quad|4\rangle\equiv\bar{X},\quad|5\rangle\equiv\bar{Y},\quad|6\rangle\equiv\bar{Z}.$ (S-72)
The zeroth step to apply our formula is to get monodromy matrices. Since the $\mathfrak{sl}_{4}\cong\mathfrak{o}_{6}$ we could use the $\mathfrak{gl}_{4}$ or $\mathfrak{o}_{6}$ formulas. Now we continue with the $\mathfrak{gl}_{4}$ convention. The $\mathfrak{gl}_{4}$ monodromy matrices can be defined in the following way. Let $e_{i}$ be the usual basis in $\mathbb{C}^{4}$. Let $e_{i,j}$ be the unit matrices in $\mathbb{C}^{4}$ as $e_{i,j}e_{k}=\delta_{j,k}e_{i}$. Let $e_{[i,j]}$ for $1\leq i<j\leq 4$ be a basis of $\mathbb{C}^{6}$ which coming from the anti-symmetrization of $\mathbb{C}^{4}\otimes\mathbb{C}^{4}$. Let $e_{[i,j],[k,l]}$ be the unit matrices in $\mathbb{C}^{6}$ as $e_{[i,j],[k,l]}e_{[a,b]}=\delta_{k,a}\delta_{l,b}e_{[i,j]}$. We can identify the six dimensional basis as
$Z$ $\equiv e_{[1,2]},$ $\bar{X}$ $\equiv e_{[2,3]},$
$Y$ $\equiv e_{[1,3]},$ $\bar{Y}$ $\equiv e_{[4,2]},$ (S-73)
$X$ $\equiv e_{[1,4]},$ $\bar{Z}$ $\equiv e_{[3,4]}.$
Let us define the 6-dimensional matrices $\mathcal{E}_{i,j}$ as
$\mathcal{E}_{i,j}=\sum_{c=1}^{4}e_{[i,c],[j,c]}.$ (S-74)
Let us define the following Lax-operators
$\mathcal{L}(u)=\sum_{i,j=1}^{4}e_{i,j}\otimes\left(\delta_{i,j}+\frac{1}{u+1/2}\mathcal{E}_{j,i}\right),\qquad\widehat{\mathcal{L}}(u)=\sum_{i,j=1}^{4}e_{i,j}\otimes\left(\delta_{i,j}+\frac{1}{-u+1/2}\mathcal{E}_{i,j}\right).$ (S-75)
The monodromy matrices
$T_{0}(u)$ $=\mathcal{L}_{0,J}(u)\dots\mathcal{L}_{0,2}(u)\mathcal{L}_{0,1}(u),$
$\widehat{T}_{0}(u)$ =^$=\widehat{\mathcal{L}}_{0,J}(u)\dots\widehat{\mathcal{L}}_{0,2}(u)\widehat{\mathcal{L}}_{0,1}(u).$ (S-76)
satisfy the $\mathfrak{gl}_{4}$ $RTT$-relations.
S-VII.1.1 $\mathfrak{o}_{4}=\mathfrak{o}_{3}\oplus\mathfrak{o}_{3}$ symmetric boundary states
Our MPS is built from a one-side operator
$\begin{split}\bar{\omega}&=\left\{S_{1},S_{2},S_{3},S_{3},S_{2},S_{1}\right\},\\ \langle\mathrm{MPS}|&=\sum_{i_{1},\dots,i_{J}}\mathrm{Tr}\left[\omega_{i_{J}}\dots\omega_{i_{1}}\right]\langle i_{1},i_{2},\dots,i_{J-1},i_{J}|.\end{split}$ (S-77)
We defined the spin operators $S_{1},S_{2},S_{3}$ which satisfy the $\mathfrak{su}_{2}$ algebra
$\left[S_{j},S_{k}\right]=\sum_{l=1}^{3}i\epsilon_{j,k,l}S_{l}.$ (S-78)
We will also need the $\mathfrak{su}_{2}$ Casimir operator
$S^{2}=S_{1}^{2}+S_{2}^{2}+S_{3}^{2}.$ (S-79)
We have finite dimensional MPS for every integer of half integer $s$ for which Casimir is $s(s+1)$.
The first step is to find the $K$-matrix. The $K$-matrix is the solution of the $KT$-relation for $J=1$
$\sum_{k=1}^{4}\sum_{a=1}^{6}\mathbf{K}_{i,k}(u)\omega_{a}\mathcal{L}_{k,j;a,b}(u)=\sum_{k=1}^{4}\sum_{a=1}^{6}\omega_{a}\widehat{\mathcal{L}}_{i,k;a,b}(-u)\mathbf{K}_{k,j}(u).$ (S-80)
The solution is
$\mathbf{K}(u)=\left(\begin{array}[]{cccc}1&\frac{2i}{2u+1}S_{1}&\frac{2i}{2u+1}S_{2}&\frac{2i}{2u+1}S_{3}\\ \frac{-2i}{2u+1}S_{1}&1&\frac{2i}{2u+1}S_{3}&\frac{-2i}{2u+1}S_{2}\\ \frac{-2i}{2u+1}S_{2}&\frac{-2i}{2u+1}S_{3}&1&\frac{2i}{2u+1}S_{1}\\ \frac{-2i}{2u+1}S_{3}&\frac{2i}{2u+1}S_{2}&\frac{-2i}{2u+1}S_{1}&1\end{array}\right).$ (S-81)
Next step is to calculate the $G$-operators which are defined in S-III.1. We need the nested $K$-matrices $\mathbf{K}^{(2)},\mathbf{K}^{(3)}$. Since $\mathbf{K}_{1,1}=1$ the second $K$-matrix is simply defined as
$\mathbf{K}_{a,b}^{(2)}(u)=\mathbf{K}_{a,b}(u)-\mathbf{K}_{a,1}(u)\mathbf{K}_{1,b}(u),$ (S-82)
for $a,b=2,3,4$ i.e. the $3\times 3$ matrix is
$\mathbf{K}^{(2)}=\left(\begin{array}[]{ccc}\mathbf{K}_{2,2}^{(2)}&\mathbf{K}_{2,3}^{(2)}&\mathbf{K}_{2,4}^{(2)}\\ \mathbf{K}_{3,2}^{(2)}&\mathbf{K}_{3,3}^{(2)}&\mathbf{K}_{3,4}^{(2)}\\ \mathbf{K}_{4,2}^{(2)}&\mathbf{K}_{4,3}^{(2)}&\mathbf{K}_{4,4}^{(2)}\end{array}\right)=\left(\begin{array}[]{ccc}1-\frac{4S_{1}^{2}}{(2u+1)^{2}}&\frac{2i}{2u+1}S_{3}-\frac{4S_{1}S_{2}}{(2u+1)^{2}}&\frac{-2i}{2u+1}S_{2}-\frac{4S_{1}S_{3}}{(2u+1)^{2}}\\ \frac{-2i}{2u+1}S_{3}-\frac{4S_{2}S_{1}}{(2u+1)^{2}}&1-\frac{4S_{2}^{2}}{(2u+1)^{2}}&\frac{2i}{2u+1}S_{1}-\frac{4S_{2}S_{3}}{(2u+1)^{2}}\\ \frac{2i}{2u+1}S_{2}-\frac{4S_{3}S_{1}}{(2u+1)^{2}}&\frac{-2i}{2u+1}S_{1}-\frac{4S_{3}S_{2}}{(2u+1)^{2}}&1-\frac{4S_{3}^{2}}{(2u+1)^{2}}\end{array}\right).$ (S-83)
The third $K$-matrix
$\mathbf{K}_{a,b}^{(3)}(u)=\mathbf{K}_{a,b}^{(2)}(u)-\mathbf{K}_{a,2}^{(2)}(u)\left[\mathbf{K}_{2,2}^{(2)}(u)\right]^{-1}\mathbf{K}_{2,b}^{(2)}(u)$ (S-84)
is more problematic because of the inverse. We can handle it with the identity
$\mathbf{K}_{2,2}^{(2)}(u+1)\mathbf{K}_{a,2}^{(2)}(u)=\left(\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)-\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\right)\mathbf{K}_{2,2}^{(2)}(u),$ (S-85)
therefore
$\mathbf{K}_{a,2}^{(2)}(u)\left[\mathbf{K}_{2,2}^{(2)}(u)\right]^{-1}=\left[\mathbf{K}_{2,2}^{(2)}(u+1)\right]^{-1}\left(\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)-\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\right).$ (S-86)
Substituting back, the third K-matrix is simplified as
$\mathbf{K}_{a,b}^{(3)}(u)=\left[\mathbf{K}_{2,2}^{(2)}(u+1)\right]^{-1}\tilde{\mathbf{K}}_{a,b}^{(3)}(u),$ (S-87)
where we introduced a $2\times 2$ matrix without inversion:
$\tilde{\mathbf{K}}_{a,b}^{(3)}(u)=\mathbf{K}_{2,2}^{(2)}(u+1)\mathbf{K}_{a,b}^{(2)}(u)-\frac{2u+3}{2u+2}\mathbf{K}_{a,2}^{(2)}(u+1)\mathbf{K}_{2,b}^{(2)}(u)+\frac{1}{2u+2}\mathbf{K}_{2,a}^{(2)}(u+1)\mathbf{K}_{2,b}^{(2)}(u).$ (S-88)
After substitution
$\tilde{\mathbf{K}}^{(3)}=\left(\begin{array}[]{cc}\tilde{\mathbf{K}}_{3,3}^{(3)}&\tilde{\mathbf{K}}_{3,4}^{(3)}\\ \tilde{\mathbf{K}}_{4,3}^{(3)}&\tilde{\mathbf{K}}_{4,4}^{(3)}\end{array}\right)=\frac{(2u+1)(2u+3)-4S^{2}}{(2u+1)(2u+3)}\left(\begin{array}[]{cc}1&\frac{2i}{2u+3}S_{1}\\ \frac{-2i}{2u+3}S_{1}&1\end{array}\right).$ (S-89)
Now we can calculate the $G$-operators
$\begin{split}\mathbf{G}^{(1)}(u)=&\mathbf{K}_{1,1}(u)=1,\\ \mathbf{G}^{(2)}(u)=&\mathbf{K}_{2,2}^{(2)}(u)=\frac{(2u+1)^{2}-4S_{1}^{2}}{(2u+1)^{2}},\\ \mathbf{G}^{(3)}(u)=&\mathbf{K}_{3,3}^{(3)}(u)=\frac{(2u+3)}{(2u+1)}\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)^{2}-4S_{1}^{2}},\\ \mathbf{G}^{(4)}(u)=&\mathbf{K}_{4,4}^{(4)}(u)=\mathbf{K}_{4,4}^{(3)}(u)-\mathbf{K}_{4,3}^{(3)}(u)\left[\mathbf{K}_{3,3}^{(3)}(u)\right]^{-1}\mathbf{K}_{3,4}^{(3)}(u)\\ =&\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)(2u+1)}.\end{split}$ (S-90)
The F-operators are
$\begin{split}\mathbf{F}^{(1)}(u)=&\left[\mathbf{G}^{(1)}(u)\right]^{-1}\mathbf{G}^{(2)}(u)=\frac{(2u+1)^{2}-4S_{1}^{2}}{(2u+1)^{2}},\\ \mathbf{F}^{(2)}(u)=&\left[\mathbf{G}^{(2)}(u)\right]^{-1}\mathbf{G}^{(3)}(u)=\frac{(2u+1)(2u+3)}{(2u+1)^{2}-4S_{1}^{2}}\frac{(2u+1)(2u+3)-4S^{2}}{(2u+3)^{2}-4S_{1}^{2}}\\ \mathbf{F}^{(3)}(u)=&\left[\mathbf{G}^{(3)}(u)\right]^{-1}\mathbf{G}^{(4)}(u)=\frac{(2u+3)^{2}-4S_{1}^{2}}{(2u+3)^{2}},\end{split},$ (S-91)
and the pseudo-vacuum ($|0\rangle=Z^{J}$) overlap is
$\mathbf{B}=S_{1}^{J}.$ (S-92)
The spectrum of the operators $S_{1},S^{2}$ can be labeled by $m=-s,-s+1,\dots,s-1,s$, therefore the eigenvalues are
$\begin{split}\mathcal{F}_{m}^{(1)}(u)&=\frac{(2u+1-2m)(2u+1+2m)}{(2u+1)^{2}},\\ \mathcal{F}_{m}^{(2)}(u)=&\frac{(2u+1)(2u+3)}{(2u+1-2m)(2u+1+2m)}\frac{(2u+1-2s)(2u+3+2s)}{(2u+3-2m)(2u+3+2m)},\\ \mathcal{F}_{m}^{(3)}(u)=&\frac{(2u+3-2m)(2u+3+2m)}{(2u+3)^{2}}.\end{split}$ (S-93)
therefore
$\begin{split}\beta_{m}&=m^{J},\\ \tilde{\mathcal{F}}_{m}^{(1)}(u)&=\frac{u^{2}+m^{2}}{u^{2}}\sqrt{\frac{u^{2}}{u^{2}+1/4}},\\ \mathcal{\tilde{\mathcal{F}}}_{m}^{(2)}(u)&=\frac{(u^{2}+1/4)(u^{2}+(s+1/2)^{2})}{(u^{2}+(m+1/2)^{2})(u^{2}+(m-1/2)^{2})}\sqrt{\frac{u^{2}}{u^{2}+1/4}},\\ \mathcal{\tilde{\mathcal{F}}}_{m}^{(3)}(u)&=\frac{u^{2}+m^{2}}{u^{2}}\sqrt{\frac{u^{2}}{u^{2}+1/4}}.\end{split}$ (S-94)
S-VII.1.2 $\mathfrak{sp}_{4}=\mathfrak{o}_{5}$ symmetric boundary states
Our MPS is built from a one-side operator
$\begin{split}\bar{\omega}&=\left\{\gamma_{5},\gamma_{1}+i\gamma_{4},\gamma_{2}+i\gamma_{3},\gamma_{2}-i\gamma_{3},\gamma_{1}-i\gamma_{4},\gamma_{5}\right\}\\ \langle\mathrm{MPS}|&=\sum_{i_{1},\dots,i_{J}}\mathrm{Tr}\left[\omega_{i_{J}}\dots\omega_{i_{1}}\right]\langle i_{1},i_{2},\dots,i_{J-1},i_{J}|.\end{split}$ (S-95)
We defined the $4\times 4$ $\gamma$-matrices which satisfy the algebra
$\left\{\gamma_{j},\gamma_{k}\right\}=2\delta_{j,k}.$ (S-96)
We can use the concrete representation
$\begin{split}\gamma_{1}&=\left(\begin{array}[]{cccc}0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&1&0&0\end{array}\right),\;\gamma_{2}=\left(\begin{array}[]{cccc}0&0&0&i\\ 0&0&i&0\\ 0&-i&0&0\\ -i&0&0&0\end{array}\right),\;\gamma_{3}=\left(\begin{array}[]{cccc}0&0&0&1\\ 0&0&-1&0\\ 0&-1&0&0\\ 1&0&0&0\end{array}\right),\\ \gamma_{4}&=\left(\begin{array}[]{cccc}0&0&i&0\\ 0&0&0&-i\\ -i&0&0&0\\ 0&i&0&0\end{array}\right),\;\gamma_{5}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\end{array}\right).\end{split}$ (S-97)
The first step is to find the $K$-matrix. The $K$-matrix is the solution of the $KT$-relation for $J=1$
$\sum_{k=1}^{4}\sum_{a=1}^{6}\mathbf{K}_{i,k}(u)\omega_{a}\mathcal{L}_{k,j;a,b}(u)=\sum_{k=1}^{4}\sum_{a=1}^{6}\omega_{a}\widehat{\mathcal{L}}_{i,k;a,b}(-u)\mathbf{K}_{k,j}(u).$ (S-98)
The solution is
$\mathbf{K}(u)=\left(\begin{array}[]{cccc}\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&a_{4}\\ 0&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&a_{3}&0\\ 0&0&0&a_{2}\end{array}\right)&\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&a_{1}\\ -a_{5}&0&0&0\\ 0&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}0&0&0&ia_{1}(u)\\ 0&0&0&0\\ 0&ia_{5}&0&0\\ 0&0&0&0\end{array}\right)\\ \left(\begin{array}[]{cccc}-1&0&0&0\\ 0&-1&0&0\\ 0&0&-a_{2}&0\\ 0&0&0&-a_{3}\end{array}\right)&\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&a_{4}&0\end{array}\right)&\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&ia_{1}&0\\ 0&0&0&0\\ ia_{5}&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}0&0&-a_{1}&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&a_{5}&0&0\end{array}\right)\\ \left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&a_{5}\\ -a_{1}&0&0&0\\ 0&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&ia_{5}&0\\ 0&0&0&0\\ ia_{1}&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}0&0&0&0\\ a_{4}&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}-a_{2}&0&0&0\\ 0&-a_{3}&0&0\\ 0&0&-1&0\\ 0&0&0&-1\end{array}\right)\\ \left(\begin{array}[]{cccc}0&0&0&ia_{5}\\ 0&0&0&0\\ 0&ia_{1}&0&0\\ 0&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}0&0&-a_{5}&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&a_{1}&0&0\end{array}\right)&\left(\begin{array}[]{cccc}a_{3}&0&0&0\\ 0&a_{2}&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right)&\left(\begin{array}[]{cccc}0&a_{4}&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)\end{array}\right),$ (S-99)
where
$a_{1}$ $=\frac{1}{u+1},$ $a_{4}$ $=\frac{-i(2u+1)}{(u-1)(u+1)},$
$a_{2}$ $=\frac{u}{u+1},$ $a_{5}$ $=\frac{(u+2)}{(u-1)(u+1)},$ (S-100)
$a_{3}$ $=\frac{u^{2}+u+1}{(u-1)(u+1)}.$
Next step is to calculate the $G$-operators which are defined in S-III.2. The first $Y^{-}(2)$ $K$-matrix
$\mathbf{k}^{(1)}(u)=\left(\begin{array}[]{cc}\mathbf{k}_{1,1}^{(1)}(u)&\mathbf{k}_{1,2}^{(1)}(u)\\ \mathbf{k}_{2,1}^{(1)}(u)&\mathbf{k}_{2,2}^{(1)}(u)\end{array}\right)=\left(\begin{array}[]{cc}\mathbf{K}_{1,1}^{(1)}(u)&\mathbf{K}_{1,2}^{(1)}(u)\\ \mathbf{K}_{2,1}^{(1)}(u)&\mathbf{K}_{2,2}^{(1)}(u)\end{array}\right),$
can be expressed with the Lax operator
$\mathbf{L}(u)=\left(\begin{array}[]{cc}\mathbf{L}_{1,1}(u)&\mathbf{L}_{1,2}(u)\\ \mathbf{L}_{2,1}(u)&\mathbf{L}_{2,2}(u)\end{array}\right)=\left(\begin{array}[]{cc}\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&\frac{u}{u-1}&0\\ 0&0&0&1\end{array}\right)&\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&\frac{-i}{u-1}&0\end{array}\right)\\ \left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&\frac{i}{u-1}\\ 0&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&\frac{u}{u-1}\end{array}\right)\end{array}\right),$ (S-101)
as
$\mathbf{k}_{a,b}^{(1)}(u)=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u)\epsilon_{c,d}\mathbf{L}_{d,b}(-u).$ (S-102)
We can define the deformed K-matrix as
$\tilde{\mathbf{k}}_{a,b}^{(1)}(u)=\mathbf{k}_{a,b}^{(1)}(u)+(u+1/2)\mathbf{L}_{1,a}(u)\mathbf{L}_{1,b}(-u).$
The first two $G$-operators are
$\begin{split}\mathbf{G}^{(1)}(u)&=\tilde{\mathbf{k}}_{1,1}^{(1)}(u)=\left(\begin{array}[]{cccc}u+1/2&0&0&0\\ 0&u+1/2&0&0\\ 0&0&\frac{u^{2}(u+1/2)}{(u-1)(u+1)}&\frac{-i(2u+1)}{(u-1)(u+1)}\\ 0&0&0&u+1/2\end{array}\right),\\ \mathbf{G}^{(2)}(u)&=\tilde{\mathbf{k}}_{2,2}^{(1)}(u)-\tilde{\mathbf{k}}_{2,1}^{(1)}(u)\left[\tilde{\mathbf{k}}_{1,1}^{(1)}(u)\right]^{-1}\tilde{\mathbf{k}}_{1,2}^{(1)}(u)=\left(\begin{array}[]{cccc}\frac{1}{u+1/2}&0&0&0\\ 0&\frac{1}{u+1/2}&0&0\\ 0&0&\frac{u(u+2)}{(u+1)^{2}(u+1/2)}&\frac{2i}{(u+1)^{2}(u+1/2)}\\ 0&0&0&\frac{1}{u+1/2}\end{array}\right).\end{split}$ (S-103)
We also need the nested $K$-matrix $\mathbf{K}^{(2)}$
$\mathbf{K}_{a,b}^{(2)}(u)=\mathbf{K}_{a,b}(u)-\sum_{\alpha,\beta=1}^{2}\mathbf{K}_{a,\alpha}(u)\widehat{\mathbf{K}}_{\alpha,\beta}^{(1)}(u)\mathbf{K}_{\beta,b}(u),$ (S-104)
for $a,b=3,4$. We need the inverse matrix
$\widehat{\mathbf{K}}^{(1)}(u)=\left(\begin{array}[]{cc}\widehat{\mathbf{K}}_{1,1}^{(1)}(u)&\widehat{\mathbf{K}}_{1,2}^{(1)}(u)\\ \widehat{\mathbf{K}}_{2,1}^{(1)}(u)&\widehat{\mathbf{K}}_{2,2}^{(1)}(u)\end{array}\right)=\left(\begin{array}[]{cc}\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&\frac{-i(2u+1)}{u(u+2)}&0\end{array}\right)&\left(\begin{array}[]{cccc}-1&0&0&0\\ 0&-1&0&0\\ 0&0&-\frac{u+1}{u}&0\\ 0&0&0&-\frac{u^{2}+u+1}{u(u+2)}\end{array}\right)\\ \left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&\frac{u^{2}+u+1}{u(u+2)}&0\\ 0&0&0&\frac{u+1}{u}\end{array}\right)&\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&\frac{-i(2u+1)}{u(u+2)}\\ 0&0&0&0\end{array}\right)\end{array}\right).$ (S-105)
After substitution we obtain the second $K$-matrix
$\mathbf{K}^{(2)}(u)=\left(\begin{array}[]{cc}\mathbf{K}_{3,3}^{(2)}(u)&\mathbf{K}_{3,4}^{(2)}(u)\\ \mathbf{K}_{4,3}^{(2)}(u)&\mathbf{K}_{4,4}^{(2)}(u)\end{array}\right)=\left(\begin{array}[]{cc}\mathbf{k}_{1,1}^{(2)}(u)&\mathbf{k}_{1,2}^{(2)}(u)\\ \mathbf{k}_{2,1}^{(2)}(u)&\mathbf{k}_{2,2}^{(2)}(u)\end{array}\right)=\\ \left(\begin{array}[]{cc}\left(\begin{array}[]{cccc}0&0&0&0\\ \frac{-i(2u+3)}{(u+1)^{2}}&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}-\frac{u}{u+1}&0&0&0\\ 0&-\frac{u^{2}+3u+3}{(u+1)^{2}}&0&0\\ 0&0&-\frac{u(u+2)}{(u+1)^{2}}&0\\ 0&0&0&-\frac{u(u+2)}{(u+1)^{2}}\end{array}\right)\\ \left(\begin{array}[]{cccc}\frac{u^{2}+3u+3}{(u+1)^{2}}&0&0&0\\ 0&\frac{u}{u+1}&0&0\\ 0&0&\frac{u(u+2)}{(u+1)^{2}}&0\\ 0&0&0&\frac{u(u+2)}{(u+1)^{2}}\end{array}\right)&\left(\begin{array}[]{cccc}0&\frac{-i(2u+3)}{(u+1)^{2}}&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)\end{array}\right).$ (S-106)
With the Lax operator
$\mathbf{L}(u)=\left(\begin{array}[]{cc}\mathbf{L}_{1,1}(u)&\mathbf{L}_{1,2}(u)\\ \mathbf{L}_{2,1}(u)&\mathbf{L}_{2,2}(u)\end{array}\right)=i\left(\begin{array}[]{cc}\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\frac{u+1}{u}&0&0\\ 0&0&\frac{u-1}{u}&0\\ 0&0&0&\frac{u-1}{u}\end{array}\right)&\left(\begin{array}[]{cccc}0&\frac{i}{u}&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)\\ \left(\begin{array}[]{cccc}0&0&0&0\\ -\frac{i}{u}&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)&\left(\begin{array}[]{cccc}\frac{u+1}{u}&0&0&0\\ 0&1&0&0\\ 0&0&\frac{u-1}{u}&0\\ 0&0&0&\frac{u-1}{u}\end{array}\right)\end{array}\right),$ (S-107)
we obtain the factorized form and the deformed matrix:
$\begin{split}\mathbf{k}_{a,b}^{(2)}(u)&=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u+1)\epsilon_{c,d}\mathbf{L}_{d,b}(-u-1),\\ \tilde{\mathbf{k}}_{a,b}^{(2)}(u)&=\mathbf{k}_{a,b}^{(2)}(u)+(u+3/2)\mathbf{L}_{1,a}(u+1)\mathbf{L}_{1,b}(-u-1).\end{split}$ (S-108)
The remaining $G$-operators are
$\begin{split}\mathbf{G}^{(3)}(u)&=\tilde{\mathbf{k}}_{1,1}^{(2)}(u)=\left(\begin{array}[]{cccc}-(u+3/2)&0&0&0\\ \frac{-i(2u+3)}{(u+1)^{2}}&-\frac{u(u+2)(u+3/2)}{(u+1)^{2}}&0&0\\ 0&0&-\frac{u(u+2)(u+3/2)}{(u+1)^{2}}&0\\ 0&0&0&-\frac{u(u+2)(u+3/2)}{(u+1)^{2}}\end{array}\right),\\ \mathbf{G}^{(4)}(u)&=\tilde{\mathbf{k}}_{2,2}^{(2)}(u)-\tilde{\mathbf{k}}_{2,1}^{(2)}(u)\left[\tilde{\mathbf{k}}_{1,1}^{(2)}(u)\right]^{-1}\tilde{\mathbf{k}}_{1,2}^{(2)}(u)=\\ &=\left(\begin{array}[]{cccc}-\frac{u(u+3)}{(u+1)(u+2)(u+3/2)}&0&0&0\\ \frac{2iu}{(u+1)^{2}(u+2)(u+3/2)}&-\frac{u(u+2)}{(u+1)^{2}(u+3/2)}&0&0\\ 0&0&-\frac{u(u+2)}{(u+1)^{2}(u+3/2)}&0\\ 0&0&0&-\frac{u(u+2)}{(u+1)^{2}(u+3/2)}\end{array}\right).\end{split}$ (S-109)
Using the basis transformation
$\mathbf{A}=\left(\begin{array}[]{cccc}0&0&0&1\\ 0&0&1&2i\\ 1&2i&0&0\\ 0&1&0&0\end{array}\right),$ (S-110)
the $B$- and $F$-operators have diagonal form and the spectrum is
$\begin{split}\mathbf{B}&=\mathbf{A}\mathrm{diag}\left((-1)^{J},(-1)^{J},1,1\right)\mathbf{A}^{-1},\\ \mathbf{F}^{(1)}(iu-1/2)&=\mathbf{A}\mathrm{diag}\left(-\frac{1}{u^{2}}\frac{(u^{2}+9/4)}{(u^{2}+1/4)},-\frac{1}{u^{2}},-\frac{1}{u^{2}},-\frac{1}{u^{2}}\right)\mathbf{A}^{-1},\\ \mathbf{F}^{(2)}(iu-2/2)&=\mathbf{A}\mathrm{diag}\left((u^{2}+1/4),(u^{2}+1/4)\frac{(u^{2}+1)}{u^{2}},(u^{2}+1/4)\frac{(u^{2}+1)}{u^{2}},(u^{2}+1/4)\right)\mathbf{A}^{-1},\\ \mathbf{F}^{(3)}(iu-3/2)&=\mathbf{A}\mathrm{diag}\left(-\frac{1}{u^{2}},-\frac{1}{u^{2}},-\frac{1}{u^{2}},-\frac{1}{u^{2}}\frac{(u^{2}+9/4)}{(u^{2}+1/4)}\right)\mathbf{A}^{-1}.\end{split}$ (S-111)
S-VII.2 Alternating $SU(4)$ spin chain
The Hamiltonian of the alternating $SU(4)$ spin chain is the following
$H=\sum_{j=1}^{2J}1-\mathbf{P}_{j,j+2}+\frac{1}{2}\mathbf{P}_{j,j+2}\mathbf{Q}_{j,j+1}+\frac{1}{2}\mathbf{Q}_{j,j+1}\mathbf{P}_{j,j+2},$ (S-112)
where
$\mathbf{P}=\sum_{i,j=1}^{4}e_{i,j}\otimes e_{j,i},\quad\mathbf{Q}=\sum_{i,j=1}^{4}e_{i,j}\otimes e_{i,j}.$ (S-113)
In the AdS/CFT context we often identify the states as the scalar fields of the ABJM as $|a\rangle\equiv Y^{a}$ for the odd sites $|a\rangle\equiv Y_{a}^{\dagger}$ for the even sites. The zeroth step to apply our formula is to get the monodromy matrix. The $\mathfrak{gl}_{4}$ monodromy matrix can be defined in the following way. Let us define the following Lax-operators
$\mathcal{L}(u)=\sum_{i,j=1}^{4}e_{i,j}\otimes\left(\delta_{i,j}+\frac{1}{u-1}e_{j,i}\right),\qquad\widehat{\mathcal{L}}(u)=\sum_{i,j=1}^{4}e_{i,j}\otimes\left(\delta_{i,j}+\frac{1}{-u-1}e_{i,j}\right).$ (S-114)
The monodromy matrix
$T_{0}(u)=\widehat{\mathcal{L}}_{0,2J}(u)\mathcal{L}_{0,2J-1}(u)\dots\widehat{\mathcal{L}}_{0,2}(u)\mathcal{L}_{0,1}(u)$ (S-115)
satisfies the $\mathfrak{gl}_{4}$ $RTT$-relation.
S-VII.2.1 $\mathfrak{gl}_{2}\oplus\mathfrak{gl}_{2}$ symmetric boundary states
Our MPS is built from a two-side operator
$\begin{split}\psi_{a,b}&=\left(\begin{array}[]{cccc}s+S_{3}+1&S_{1}+iS_{2}&0&0\\ S_{1}-iS_{2}&s-S_{3}+1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)_{a,b},\\ \langle\mathrm{MPS}|&=\sum_{i_{1},\dots,i_{J}}\mathrm{Tr}\left[\psi_{i_{2J-1},i_{2J}}\dots\psi_{i_{1},i_{2}}\right]\langle i_{1},i_{2},\dots,i_{2J-1},i_{2J}|,\end{split}$ (S-116)
where $S_{1},S_{2},S_{3}$ are the spin operators for the irreducible spin $s$ representation.
The first step is to find the $K$-matrix. The $K$-matrix is the solution of the $KT$-relation for $J=1$
$\sum_{a,b,k,l=1}^{4}\mathbf{K}_{i,k}(u)\psi_{a,b}\widehat{\mathcal{L}}_{k,l;b,d}(u)\mathcal{L}_{l,j;a,c}(u)=\sum_{a,b,k,l=1}^{4}\psi_{a,b}\widehat{\mathcal{L}}_{i,k;b,d}(-u)\mathcal{L}_{k,l;a,c}(-u)\mathbf{K}_{l,j}(u).$ (S-117)
The solution is
$\mathbf{K}(u)=\left(\begin{array}[]{cccc}a(u)+uS_{3}&uS_{-}&0&0\\ uS_{+}&a(u)-uS_{3}&0&0\\ 0&0&b(u)&0\\ 0&0&0&b(u)\end{array}\right),$ (S-118)
where $S_{\pm}=S_{1}\pm iS_{2}$ and
$\begin{split}a(u)&=-\frac{1}{2}u^{2}+\frac{1}{2}u+s+1,\\ b(u)&=(u-1)(\frac{1}{2}u-s-1).\end{split}$ (S-119)
Next step is to calculate the $G$-operators which are defined in S-III.3. We can see that the $\mathbf{K}_{4,1}$ component is not invertible therefore we have to rotate the boundary state and the $K$-matrix. We use the $GL(4)$ transformation
$g=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&-\frac{\mathfrak{b}_{2}}{2}&1&0\\ -\frac{\mathfrak{b}_{1}}{2}&0&0&1\end{array}\right).$ (S-120)
The co-product is
$\Delta(g)=\left(g\otimes\left(g^{-1}\right)^{t}\right)^{\otimes J},$ (S-121)
for which
$\left[\Delta(g),H\right]=0.$ (S-122)
The $K$-matrix and the boundary state transforms as
$\begin{split}\psi_{a,b}^{g}&=\sum_{c,d=1}^{4}\psi_{c,d}g_{c,a}\left(g^{-1}\right)_{b,d},\\ \mathbf{K}^{g}(u)&=\left(g^{-1}\otimes\mathbf{1}\right)\mathbf{K}(u)\left(g\otimes\mathbf{1}\right).\end{split}$ (S-123)
In the following we work with the transformed quantities and take the $\mathfrak{b}_{1},\mathfrak{b}_{2}\to 0$ limit only for the final formula.
The pseudo-vacuum is $|0\rangle=\left(|1\rangle\otimes|4\rangle\right)^{\otimes J}$ and the pseudo-vacuum overlap is
$\mathbf{B}=\langle\Psi^{g}|0\rangle=(\psi_{1,4}^{g})^{J}=\mathfrak{b}_{1}^{J}\left(\frac{S_{3}+s+1}{2}\right)^{J}.$ (S-124)
We need the nested $K$-matrix
$\mathbf{K}_{a,b}^{(2)}(u)=\mathbf{K}_{a,b}^{g}(u)-\mathbf{K}_{a,1}^{g}(u)\left[\mathbf{K}_{4,1}^{g}(u)\right]^{-1}\mathbf{K}_{4,b}^{g}(u),$ (S-125)
for $a,b=2,3$. We can handle the inverse with the identity
$\mathbf{K}_{4,1}^{g}(u+1)\mathbf{K}_{a,1}^{g}(u)=\mathbf{K}_{a,1}^{g}(u+1)\mathbf{K}_{4,1}^{g}(u),$ (S-126)
for $a=2,3$, therefore
$\mathbf{K}_{a,1}^{g}(u)\left[\mathbf{K}_{4,1}^{g}(u)\right]^{-1}=\left[\mathbf{K}_{4,1}^{g}(u+1)\right]^{-1}\mathbf{K}_{a,1}^{g}(u+1).$ (S-127)
Substituting back, the nested K-matrix is simplified as
$\mathbf{K}_{a,b}^{(2)}(u)=\left[\mathbf{K}_{4,1}^{g}(u+1)\right]^{-1}\tilde{\mathbf{K}}_{a,b}^{(2)}(u),$ (S-128)
where we introduced a $2\times 2$ matrix without inversion:
$\tilde{\mathbf{K}}_{a,b}^{(2)}(u)=\mathbf{K}_{4,1}^{g}(u+1)\mathbf{K}_{a,b}^{g}(u)-\mathbf{K}_{a,1}^{g}(u+1)\mathbf{K}_{4,b}^{g}(u).$ (S-129)
After substitution
$\tilde{\mathbf{K}}^{(2)}=\left(\begin{array}[]{cc}\tilde{\mathbf{K}}_{2,2}^{(2)}&\tilde{\mathbf{K}}_{2,3}^{(2)}\\ \tilde{\mathbf{K}}_{3,2}^{(2)}&\tilde{\mathbf{K}}_{3,3}^{(2)}\end{array}\right)=\frac{\mathfrak{b}_{1}\mathfrak{b}_{2}}{4}u(u^{2}-1)(u-2s-2)\left(\begin{array}[]{cc}\frac{1}{\mathfrak{b}_{2}}\frac{S_{3}+s+1}{u}+\frac{1}{\mathfrak{b}_{2}}&0\\ 1&\frac{1}{\mathfrak{b}_{2}}\frac{S_{3}+s+1}{u}-\frac{1}{\mathfrak{b}_{2}}\end{array}\right).$ (S-130)
Now we can calculate the $G$-operators
$\begin{split}\mathbf{G}^{(1)}(u)=&\mathbf{K}_{4,1}^{g}(u)=-\frac{\mathfrak{b}_{1}}{2}u(u-S_{3}-s-2),\\ \mathbf{G}^{(2)}(u)=&\mathbf{K}_{3,2}^{(2)}(u)=-\frac{\mathfrak{b}_{2}}{2}\frac{u(u-1)(u-2s-2)}{(u-S_{3}-s-1)},\\ \mathbf{G}^{(3)}(u)=&\mathbf{K}_{2,3}^{(2)}(u)-\mathbf{K}_{2,2}^{(2)}(u)\left[\mathbf{K}_{3,2}^{(2)}(u)\right]^{-1}\mathbf{K}_{3,3}^{(2)}(u)=\\ =-&\frac{1}{2\mathfrak{b}_{2}}\frac{(u-1)(u-2s-2)(u+S_{3}+s+1)}{u}.\end{split}$ (S-131)
The F-operators are
$\begin{split}\mathbf{F}^{(1)}(u)=&\left[\mathbf{G}^{(1)}(u)\right]^{-1}\mathbf{G}^{(2)}(u)=\frac{\mathfrak{b}_{2}}{\mathfrak{b}_{1}}\frac{(u-1)(u-2s-2)}{(u-S_{3}-s-1)(u-S_{3}-s-2)},\\ \mathbf{F}^{(2)}(u)=&\left[\mathbf{G}^{(2)}(u)\right]^{-1}\mathbf{G}^{(3)}(u)=\frac{1}{\mathfrak{b}_{2}^{2}}\frac{(u+S_{3}+s+1)(u-S_{3}-s-1)}{u^{2}}.\end{split}$ (S-132)
The spectrum of the operators $S_{3}$ can be labeled by $m=-s,-s+1,\dots,s-1,s$, therefore
$\begin{split}\beta_{m}&=\mathfrak{b}_{1}^{J}\left(\frac{s+m+1}{2}\right)^{J},\\ \tilde{\mathcal{F}}_{m}^{(1)}(u)&=\frac{\mathfrak{b}_{2}}{\mathfrak{b}_{1}}\frac{(u+i/2)(u+2is+3i/2)}{(u+im+is+i/2)(u+im+is+3i/2)},\\ \mathcal{\tilde{\mathcal{F}}}_{m}^{(2)}(u)&=\frac{1}{\mathfrak{b}_{2}^{2}}\frac{u^{2}+(m+s+1)^{2}}{u^{2}}\sqrt{\frac{u^{2}}{u^{2}+1/4}}.\end{split}$ (S-133)
The normalized overlap can be written as
$\frac{\langle\mathrm{MPS}_{g}|\bar{u}\rangle}{\sqrt{\langle\bar{u}|\bar{u}\rangle}}=\mathfrak{b}_{1}^{J-r_{1}}\mathfrak{b}_{2}^{r_{1}-r_{2}}\sqrt{\frac{\det G^{+}}{\det G^{-}}}\times\\ \left[\sum_{m=-s}^{s}\left(\frac{s+m+1}{2}\right)^{J}\prod_{j=1}^{r_{1}}\frac{(u_{j}^{1}+i/2)(u_{j}^{1}+2is+3i/2)}{(u_{j}^{1}+im+is+i/2)(u_{j}^{1}+im+is+3i/2)}\prod_{j=1}^{r_{2}/2}\frac{(u_{j}^{2})^{2}+(m+s+1)^{2}}{u_{j}^{2}\sqrt{(u_{j}^{2})^{2}+1/4}}\right].$ (S-134)
In the $\mathfrak{b}_{1},\mathfrak{b}_{2}\to 0$ limit the overlap is non-vanishing only when $J=r_{1}=r_{2}$.
S-VIII Overlaps with odd number of chiral Bethe roots
So far, for simpler presentation, we have focused on Bethe vectors where each chiral nodus has an even number of Bethe roots. Small changes are sufficient for generalization. In the following, we show how the formulas are modified in the presence of odd numbers of chiral Bethe roots.
When a chiral node $s$ have odd number of Bethe roots we define the quantum numbers $r_{s}^{+}=\frac{r_{s}-1}{2}$ and $r_{s}^{0}=1$. We also introduce the notations $u_{k}^{+,s}\equiv u_{k}^{s}$, $u_{k}^{-,s}\equiv u_{r_{s}+1-k}^{s}$ for $k=1,\dots,r_{s}^{+}$ and $u^{0,s}\equiv u_{r_{s}^{+}+1}^{s}$. In the pair structure limit $u_{k}^{-,s}=-u_{k}^{+,s}$ for $k=1,\dots,r_{s}^{+}$ and $u^{0,s}=0$. When a node $s$ is achiral or chiral with even number of Bethe roots then $r_{s}^{0}=0$. We also define the following sets $\bar{u}^{\pm,s}=\left\{u_{k}^{\pm,s}\right\}_{k=1}^{r_{s}^{+}}$, $\bar{u}^{0,s}=\left\{u^{0,s}\right\}_{k=1}^{r_{s}^{0}}$ and $\bar{u}^{\pm}=\left\{\bar{u}^{\pm,1},\bar{u}^{\pm,2},\dots,\bar{u}^{\pm,n_{+}}\right\}$, $\bar{u}^{0}=\left\{\bar{u}^{0,1},\bar{u}^{0,2},\dots,\bar{u}^{0,n_{+}}\right\}$.
The expressions $\phi_{k}^{s}$ can be grouped according to the pair structure in the same way as the Bethe roots. According to the decompositions $\bar{u}=\bar{u}^{+}\cup\bar{u}^{0}\cup\bar{u}^{-}$ and $\bar{\phi}=\bar{\phi}^{+}\cup\bar{\phi}^{0}\cup\bar{\phi}^{-}$ the Gaudin matrix has the following block form
$G=\left(\begin{array}[]{ccc}A^{++}&A^{+0}&A^{+-}\\ A^{0+}&A^{00}&A^{0-}\\ A^{-+}&A^{-0}&A^{--}\end{array}\right).$ (S-135)
The Gaudin determinants are
$\det G^{+}=\left|\begin{array}[]{cc}A^{++}+A^{+-}&A^{+0}\\ A^{0+}&\frac{1}{2}A^{00}\end{array}\right|,\quad\det G^{-}=2\left|A^{++}-A^{+-}\right|.$ (S-136)
The overlap function is only minimally modified:
$\frac{\langle\mathrm{MPS}|\bar{u}\rangle}{\sqrt{\langle\bar{u}|\bar{u}\rangle}}=\left[\sum_{k=1}^{d_{B}}\beta_{k}\prod_{j=1}^{n_{+}}\left(\gamma_{k}^{(j)}\right)^{r_{j}^{0}}\tilde{\mathcal{F}}_{k}^{(j)}(\bar{u}^{+,j})\right]\sqrt{\frac{\det G^{+}}{\det G^{-}}}.$ (S-137)
We introduced a set of new parameters $\gamma_{k}^{(j)}\in\mathbb{C}$ for $k=1.\dots,d_{B}$ and chiral nodes $j$. These parameters can be easily determined from numerical calculations.